### v3

# A Lie algebraic pattern behind logarithmic CFTs

Abstract.
We introduce a purely Lie algebraic formalization of the Feigin–Tipunin’s geometric construction of logarithmic CFTs/VOAs. After reformulating the geometric representation theory of FT construction under this new setting, within this framework, we uniformly construct the (multiplet) principal W-algebras at positive integer level associated with any simple Lie algebra $\mathfrak{g}$ and Lie superalgebra $\mathfrak{osp}(1|2r)$, thereby establishing Weyl-type character formulas and simplicity theorems that extend the second author’s previous results.
1. Introduction
The study of logarithmic conformal field theory (LCFT) has become increasingly important in recent years. However, LCFT is much more complicated than rational cases and has not been well studied. In 2010, Feigin–Tipunin [27] proposed a geometric construction of LCFTs called the Feigin–Tipunin (FT) construction, i.e., the 0-th sheaf cohomology $H^{0}(G\times_{B}V)$ of a $G$-equivariant VOA-bundle $G\times_{B}V$ over the flag variety $G/B$, and claimed that multiplet W-algebras [^1] In the case of $\mathfrak{g}=\mathfrak{sl}_{2}$ with the principal nilpotent element, this is one of the most famous and well studied LCFTs called triplet Virasoro algebra (see e.g. [7]). Here the “multiplet W-(super)algebra” is intended to be a generic term for variations of the triplet Virasoro algebra (e.g., the cases of singlet/doublet, higher rank, general W-algebra, etc.). defined by the joint-kernel of screening operators can be constructed and studied by their method. The second author [45, 46] gave rigorous proofs of the claims in [27] and several fundamental properties on the multiplet W-algebra and the corresponding principal W-algebra for simply-laced cases. In [14], he also studied the $V^{(p)}$-algebra (== doublet affine $\mathfrak{sl}_{2}$) [1, 4] by combining the method in [45, 46] with the inverse quantum Hamiltonian reduction [2]. On the other hand, a close examination of [45, 46] suggests that, although formulated in the framework of VOAs, the essential arguments ultimately reduce to properties of an underlying Lie algebraic setting. As detailed below, extracting such Lie algebraic setting and argument not only provides a clear perspective for the studies of LCFTs, but also sheds light [47, 48] on the expected correspondence between LCFTs and $3$-manifolds via the $\hat{Z}-$invariants [15, 16, 17, 30].
In Part I, we will introduce such a purely Lie algebraic setting, tentatively named shift system, and prove that the main results of [45] can be understood and derived from the setting (Theorem 1.1). In other words, if a VOA-module $V$ (e.g. irreducible lattice VOA-module) fits into a shift system, the corresponding multiplet W-algebra satisfies Theorem 1.1. In particular, as it has the FT construction $H^{0}(G\times_{B}V)$, we can study it using the geometric representation theory. As demonstrated in Section 2.4, the verification of the axiom of shift system is basically reduced to a relatively easy computation (== Serre relation of long screening operators) on $V$ and the rank $1$ case (== “Felder complex” of short screening operators). Hence, the shift system provides a useful framework for reducing the study of higher rank LCFTs to much easier cases: the rank $1$ cases and free field algebras. Moreover, with some information about the corresponding “W-algebra” $H^{0}(G\times_{B}V)^{G}$ (e.g. Kazhdan–Lusztig decomposition), the main results of [46] are proved (Section 2.2). Using these results, in Part II, we will give the shift system and the FT constructions corresponding to the (multiplet) principal W-(super)algebras $\mathbf{W}^{k}(\mathfrak{g})$ for any simple Lie algebra $\mathfrak{g}$ and Lie superalgebra $\mathfrak{g}=\mathfrak{osp}(1|2n)$, respectively, and prove almost all the main results of [45, 46] for these new cases (Therem 1.2). At first glance, these results themselves may appear to be just a straightforward extension of [45, 46]. However, as mentioned briefly at the end of the last paragraph, the value of this paper lies in the change of perspective, which greatly increases the transparency and flexibility of the discussion in LCFTs, by focusing not on individual examples of LCFT and their complicated VOA-structures but on a simple Lie algebraic pattern behind them in common. In that sense, the application to the two examples in Part II is a demonstration of this new perspective which will be further developed itself and applied to much broader class of expected LCFTs corresponding to $3$-manifolds in the future (see the last of introduction below).
Let us describe the overview of this paper. For a finite dimensional simple Lie algebra $\mathfrak{g}$ of rank $r$, we consider a triple $(\Lambda,\uparrow,\{V_{\lambda}\}_{\lambda\in\Lambda})$, named shift system (Definition 1.6). Here $\Lambda$ is a module over the Weyl group $W$ of $\mathfrak{g}$, $\uparrow\colon W\times\Lambda\rightarrow\mathfrak{h}^{\ast}$ is a map (shift map) satisfying a few simple axioms, and $V_{\lambda}$ is a graded weight $B$-module with a parameter $\lambda\in\Lambda$ and a “Felder complex” [21] for each direction $i\in I:=\{1,\dots,r\}$. Recall that in [27, 45, 46], $V_{\lambda}$ is an irreducible module $V_{\sqrt{p}(Q+\lambda)}$ over the lattice VOA $V_{\sqrt{p}Q}$ with a conformal grading $V_{\lambda}=\bigoplus_{\Delta}V_{\lambda,\Delta}$ and a parameter $\lambda\in\Lambda\simeq(\sqrt{p}Q)^{\ast}/\sqrt{p}Q\simeq\tfrac{1}{p}Q^{\ast}/Q$, and the long/short screening operators define $B$- and $W$-actions on $V_{\lambda}$ and $\Lambda$, respectively. The shift of Cartan weights by the short screening operators satisfies the axioms of the shift map above. The situation in [14] is a little more complicated, but essentially the same. Therefore, the shift system can be regarded as an abstraction of the relationship between such VOA-modules and screening operators acting on them, forgetting the VOA-structure. The first main theorem of Part I asserts that all the main results of [45] hold even under such abstraction.
Theorem 1.1.
Let $(\Lambda,\uparrow,\{V_{\lambda}\}_{\lambda\in\Lambda})$ is a shift system (see Definition 1.6).
  1. (1)
    (Feigin–Tipunin construction) The evaluation map
    $\operatorname{ev}\colon H^{0}(G\times_{B}V_{\lambda})\rightarrow\bigcap_{i\in I}\ker Q_{i,\lambda}|_{V_{\lambda}},\ \ s\mapsto s(\operatorname{id}_{G/B})$
    is injective, and is isomorphic iff $\lambda\in\Lambda$ satisfies the following “weak condition”:
    (weak) For any $(i,j)\in I\times I$, $\lambda\in\Lambda^{\sigma_{j}}$ or $({\sigma_{j}}\uparrow{\lambda},\alpha_{i}^{\vee})=-\delta_{ij}$.
    Note that the image of $H^{0}(G\times_{B}V_{\lambda})$ is the maximal $G$-submodule [^2] Namely, the maximal $B$-submodule such that its $B$-action can be extended to the $G$-action. of $V_{\lambda}$.
  2. (2)
    (Borel–Weil–Bott-type theorem) For a minimal expression $w_{0}=\sigma_{i_{l(w_{0})}}\cdots\sigma_{i_{1}}\sigma_{i_{0}}$ of the longest element $w_{0}$ of $W$ (where $\sigma_{i_{0}}=\operatorname{id}$ for convenience), if $\lambda\in\Lambda$ satisfies the following “strong condition”:
    (strong) For any $0\leq m\leq N-1$, $(\sigma_{i_{m}}\cdots\sigma_{i_{0}}\uparrow\lambda,\alpha_{i_{m+1}}^{\vee})=0$,
    then we have a natural $G$-module isomorphism
    $H^{n}(G\times_{B}V_{\lambda})\simeq H^{n+l(w_{0})}(G\times_{B}V_{w_{0}\ast\lambda}(w_{0}\uparrow\lambda)).$
Note that the weak/strong conditions in Theorem 1.1 are special cases of Definition 1.6(2b).
Since the constructions and homomorphisms in Theorem 1.1 are natural, they are compatible with additional algebraic (e.g, VOA-module) structures. In particular, if $V_{0}$ is a VOA and $V_{\lambda}$ are $V_{0}$-modules, then $H^{0}(G\times_{B}V_{0})$ has the induced VOA-structure and $H^{n}(G\times_{B}V_{\lambda})$ are $H^{0}(G\times_{B}V_{0})$-modules (see [46, Section 2.1, Corollary 2.21]). It allows us to develop a geometric representation theory of multiplet W-(super)algebras, traditionally defined in the form of the right-hand side of Theorem 1.1(1). Except for rank 1 cases, these VOAs are difficult to study algebraically due to their complicated VOA structures. However, the geometric construction allows us to avoid the bottleneck and obtain a variety of results qualitatively, as in [45, 46, 14].
In Part II, we demonstrate use and utility of the shift system and Theorem 1.1 by applying them to specific VOSAs. Let us consider the rescaled root lattice $\sqrt{p}Q$ for some $p\in\mathbb{Z}_{\geq 1}$. We consider the case where $\sqrt{p}Q$ is positive-definite integral, namely, $p\in r^{\vee}\mathbb{Z}_{\geq 1}$ for the lacing number $r^{\vee}$ or $p$ is odd and $\mathfrak{g}=B_{r}$. In the first case, $\sqrt{p}Q$ is even and we consider $V_{\lambda}=V_{\sqrt{p}(Q+\lambda)}$. On the other hand, in the second case, $\sqrt{p}Q$ is odd and we need a modification $V_{\lambda}=V_{\sqrt{p}(Q+\lambda)}\otimes F$ by the free fermion $F$. In Section 2.4, from minimal natural assumptions (2.8) or (2.9), shift systems $(\Lambda,\uparrow,\{V_{\lambda}\}_{\lambda\in\Lambda})$ are uniquely constructed (Theorem 2.13). As noted in the second paragraph above, only relatively easy computations on $V_{\lambda}$ and some basic results in rank $1$ cases ([7] and [9, 8]) are used here. Therefore, Theorem 1.1 is applied to our cases (for more detail, see Section 3).
Let us explain the two consequences of Theorem 1.1(2), namely, Weyl-type character formula and Simplicity theorem (see Section 1.5 and 2.2). The natural $G$-action on $H^{0}(G\times_{B}V_{\lambda})$ gives the decomposition
(1.1) $H^{0}(G\times_{B}V_{\lambda})\simeq\bigoplus_{\beta\in P_{+}}L_{\beta}\otimes\mathcal{W}_{-\beta+\lambda},$
where $L_{\beta}$ is the irreducible $\mathfrak{g}$-module with highest weight $\beta$, and $\mathcal{W}_{-\beta+\lambda}$ is the multiplicity of a weight vector of $L_{\beta}$. In Part II, we consider additional VOA structures, where $\mathcal{W}_{0}\simeq H^{0}(G\times_{B}V_{0})^{G}$ is a subalgebra of the multiplet W-(super)algebra $H^{0}(G\times_{B}V_{0})$ and $\mathcal{W}_{-\beta+\lambda}$ is a $\mathcal{W}_{0}$-modules. If $\lambda$ satisfies (strong), then by combining $H^{n>0}(G\times_{B}V_{\lambda})\simeq 0$ with the Atiyah–Bott localization formula, we obtain the Weyl-type character formula
(1.2) $\operatorname{ch}_{q}H^{0}(G\times_{B}V_{\lambda})=\sum_{\beta\in P_{+}}\operatorname{dim}L_{\beta}\operatorname{ch}_{q}\mathcal{W}_{-\beta+\lambda}=\sum_{\beta\in P_{+}}\operatorname{dim}L_{\beta}\sum_{\sigma\in W}(-1)^{l(\sigma)}\operatorname{ch}_{q}V_{\sigma\ast\lambda}^{h=\beta-\sigma\uparrow\lambda}.$
On the other hand, by combining the case $n=0$ with Serre duality, we have $H^{0}(G\times_{B}V_{\lambda})\simeq H^{0}(G\times_{B}V_{\lambda^{\prime}})^{\ast}$ for some $\lambda^{\prime}\in\Lambda$. In the VOA side, it leads the self-duality and simplicity of the vacuum case $H^{0}(G\times_{B}V_{0})$. By the quantum Galois theory [19, 43], $\mathcal{W}_{-\beta}$ are also simple as $\mathcal{W}_{0}$-modules. In addition, if a Kazhdan–Lusztig-type character formula for the simple quotient of $\mathcal{W}_{-\beta+\lambda}$ is known, by comparing it with (1.2), we can extend the simplicity to the whole $\lambda\in\Lambda$ satisfying (strong). For more detail, see Section 2.2.
Let us go back to our special cases above. In our cases, $\mathcal{W}_{0}$ is the principal W-(super)algebra $\mathbf{W}^{k}(\mathfrak{g})$ or $\mathbf{W}^{k}(\mathfrak{osp}(1|2r))$, and thus we have the Kazhdan–Lusztig type character formula derived from [37, 38] and the exactness of ++-reduction $H^{0}_{\mathrm{DS},+}(\cdot)$ in [10] (however, in the second case we assume the latter). By applying the discussion in the last paragraph, we can extend the main results of [5] to our cases (Theorem 4.3 and 4.8).
Theorem 1.2.
Let us consider the setup in Theorem 2.13 and Section 4.1.
Then for each case, we have $\mathcal{W}_{0}\simeq\mathbf{W}^{k}(\mathfrak{g})\simeq\mathbf{W}^{\check{k}}({}^{L}\mathfrak{g})$ and $\mathcal{W}_{0}\simeq\mathbf{W}^{k}(\mathfrak{osp}(1|2r))\simeq\mathbf{W}^{\check{k}}(\mathfrak{osp}(1|2r))$, respectively. Furthermore, in the first case, for any $\alpha\in P_{+}\cap Q$ and $\lambda\in\Lambda$ such that $(p\lambda_{\bullet}+\rho^{\vee},{}^{L}\theta)\leq p$, $\mathcal{W}_{-\alpha+\lambda}$ and $H^{0}(G\times_{B}V_{\lambda})$ are simple as $\mathcal{W}_{0}$- and $H^{0}(G\times_{B}V_{0})$-modules, respectively. In the second case, under similar conditions with the restriction $\lambda^{\bullet}=0$ and the assumption that the ++-reduction is exact, the same simplicity theorem holds. In the first case, the above simplicity of $\mathcal{W}_{-\alpha+\lambda}$ also leads to the simplicity and duality of the Arakawa–Frenkel modules $\check{\mathbf{T}}^{m}_{p\lambda_{\bullet},\alpha+\lambda^{\bullet}}\simeq\mathbf{T}^{1/p}_{\alpha+\lambda^{\bullet},p\lambda_{\bullet}}$.
The decomposition (1.1) with Theorem 1.2 is regarded as the Schur-Weyl type duality for the multiplet principal W-algebras. In the non-simply laced case, in light of [43], the tensor category of $\mathbf{W}^{k}(\mathfrak{g})$-modules is expected to exhibit a symmetry governed by the Langlands dual Lie algebra $^{L}\mathfrak{g}$.
Finally, let us discuss two directions to which Theorem 1.1 points. Our geometric representation theory of FT construction has similar aspects to the modular representation theory. Indeed, in our special cases, the strong condition in Lemma 2.9(1) also appear in the BWB theorem for $^{L}G$ [32, II.5]. Ultimately, such a similarity should come from the log Kazhdan–Lusztig correspondence (e.g. [24, 27, 44, 42, 29, 13]): an expected categorical equivalence between a multiplet $W$-algebra and a “quantum group”. For example, in our case, $H^{0}(G\times_{B}V_{\sqrt{p}Q})$ is expected to correspond to the small quantum group $u_{\zeta}({}^{L}\mathfrak{g})$ at $\zeta=\frac{\pi i}{p}$. [^3] On the other hand, our super case is expected to correspond to some small quantum group at $\zeta=\tfrac{2\pi i}{2m-1}$ (see [9]). Although the case $\mathfrak{g}=\mathfrak{sl}_{2}$ was already treated in [29], it seems necessary to develop both our geometric methods and W-algebraic approaches for higher rank generalizations (see Remark 4.4).
On the other hand, under the correspondence of the BWB-type theorems above, Theorem 1.1(1) clearly does not hold in the modular representation theory [32]. In this regard, the second author conceived the following idea a few years ago: Let $\vec{\lambda}=(\lambda_{1},\dots,\lambda_{N})\in\Lambda_{1}\times\cdots\times\Lambda_{N}$ be a parameter sequence and $\tilde{V}_{\vec{\lambda}}$ a shift system with respect to $\lambda_{1}\in\Lambda_{1}$. If we consider the situation such that for each $0\leq n\leq N-1$, an appropriate quotient
$\underbrace{\tilde{H}^{0}(G\times_{B}\tilde{H}^{0}(G\times_{B}\cdots\tilde{H}^{0}(G\times_{B}}_{\text{$n$-times}}\tilde{V}_{\vec{\lambda}})\cdots))\ \ \text{of}\ \ \underbrace{{H}^{0}(G\times_{B}\tilde{H}^{0}(G\times_{B}\cdots\tilde{H}^{0}(G\times_{B}}_{\text{$n$-times}}\tilde{V}_{\vec{\lambda}})\cdots))$
is a shift system with respect to $\lambda_{n+1}\in\Lambda_{n+1}$, then we finally obtain the nested FT construction
$\underbrace{H^{0}(G\times_{B}\tilde{H}^{0}(G\times_{B}\cdots\tilde{H}^{0}(G\times_{B}}_{\text{$N$-times}}\tilde{V}_{\vec{\lambda}})\cdots)).$
Because a shift system appears at each stage of nesting, Theorem 1.1 can be applied repeatedly. In particular, repeated application of (1.2)444 To compute $\operatorname{ch}_{q}\tilde{H}^{0}(G\times_{B}\cdots)$ from $\operatorname{ch}_{q}H^{0}(G\times_{B}\cdots)$ using (1.2), we assume a suitable match between them. in the case $\mathfrak{g}=\mathfrak{sl}_{2}$, surprisingly, almost coincides with the $\hat{Z}-$invariant [30] of the negative definite Seifert $3$-manifold! This suggests that such nested FT constructions yield rich examples of LCFT whose representation theories are to some extent reducible and controllable by the geometric representation theory of FT construction, providing a blueprint for the “dictionary” between LCFTs and negative definite plumbed $3$-manifolds expected in [15, 16, 17], etc. For $\mathfrak{g}=\mathfrak{sl}_{2}$, the second author already achieved the nested FT construction and reconstructed the $\hat{Z}-$invariants of these $3$-manifolds [47, 48].
Acknowledgments: H.L. thanks to Antun Milas for motivating questions, and to Drazen Adamovic and Naoki Genra for useful discussions. S.S. thanks to Naoki Genra and Hiroshi Yamauchi for useful comments on naming. The authors thank to Tomoyuki Arakawa for useful feedback. The majority of this work was done when the authors were postdocs in YMSC, Tsinghua University.
Part I Shift system
In Part I, we describe the discussion of [27, 45, 46] in an axiomatic (in other words, purely Lie algebraic) manner. We also generalize the discussion in these papers to include non-simply laced cases.
1.1. Preliminary from Lie algebra
We present the basic notations and facts used throughout this paper. Let $\mathfrak{g}=\mathfrak{n}_{-}\oplus\mathfrak{h}\oplus\mathfrak{n}_{+}$ be a finite dimensional simple Lie algebra of rank $r$ and lacing number $r^{\vee}$, and its triangular decomposition. Let $\Delta=\Delta_{+}\sqcup\Delta_{-}$ be the root system of $\mathfrak{g}$ (resp. positive roots and negative roots). For the Weyl group $W$ of $\mathfrak{g}$, we consider the canonically normalized $W$-invariant bilinear form $(\cdot,\cdot)=(\cdot,\cdot)_{Q}$, namely, for a long root $\alpha\in\Delta^{l}$, the length $|\alpha|^{2}$ is always $2$ (denote $\Delta^{s}$ the short roots). For a root $\alpha\in\Delta$, the coroot $\alpha^{\vee}$ is defined by $\tfrac{2}{|\alpha|^{2}}\alpha$. For $1\leq i\leq r$, $\alpha_{i}$, $\alpha_{i}^{\vee}$, $\alpha_{i}^{\ast}$, $\varpi_{i}$ and $\sigma_{i}\in W$ are the simple root, simple coroot, fundamental coweight (dual of $\alpha_{i}$), fundamental weight (dual of $\alpha_{i}^{\vee}$), and the simple reflection $\sigma_{i}(\mu)=\mu-(\mu,\alpha_{i}^{\vee})\alpha_{i}$ corresponding to $\alpha_{i}$, respectively. In this paper, the labeling of the Dynkin diagrams of $\mathfrak{g}$ are given by
respectively. We sometimes identify the set of simple roots $\Pi=\{\alpha_{1},\dots,\alpha_{r}\}$ with the set $I=\{1,\dots,r\}$. The fundamental weight $\varpi_{i}$ is the dual vector of the coroot. Denote $Q=\sum_{i\in I}\mathbb{Z}\alpha_{i}$, $Q^{\vee}=\sum_{i\in I}\mathbb{Z}\alpha_{i}^{\vee}$, $Q^{\ast}=\check{P}=\sum_{i\in I}\mathbb{Z}\alpha_{i}^{\ast}$ and $P=\sum_{i\in I}\mathbb{Z}\varpi_{i}$ the root lattice, coroot lattice, coweight lattice and weight lattice, respectively. The set of dominant integral weights $P_{+}$ and dominant integral coweights $\check{P}_{+}=Q^{\ast}_{+}$ are given by $P_{+}=\sum_{i\in I}\mathbb{Z}_{\geq 0}\varpi_{i}$, $\check{P}=\sum_{i\in I}\mathbb{Z}_{\geq 0}\alpha_{i}^{\ast}$, respectively. We use the letter $P_{\mathrm{min}}\subseteq P_{+}$ for the family of minuscule weights. Let $G$ be the corresponding simply-connected simple algebraic group with $B$ the Borel subgroup. Given a subset $J\subset I$, we denote by $P_{J}$ the corresponding parabolic subgroup [^5] Throughout the paper, we use the negative parabolic/Borel subalgebras. and $j\in J$, by $SL_{2}^{j}$ the corresponding subgroup isomorphic to $SL_{2}$. Denote $\rho=\sum_{i\in I}\varpi_{i}$ and $\rho^{\vee}=\sum_{i\in I}\alpha_{i}^{\ast}$ the Weyl vector and Weyl covector, $\theta$ and $\theta_{s}$ the highest (long) root and highest short root, $h$ and $h^{\vee}$ the Coxeter and dual Coxeter number, respectively. For $\sigma\in W$, $l(\sigma)$ and $w_{0}$ denote the length of $\sigma$ and the longest element in $W$, respectively. For a minimal expression $\sigma=\sigma_{i_{n}}\cdots\sigma_{i_{1}}$ of $\sigma\in W$, we sometimes use $\sigma_{i_{0}}=\operatorname{id}$ for convenience. For $\beta\in P_{+}$, denote $L_{\beta}$ the finite-dimensional irreducible $\mathfrak{g}$-module with the highest weight $\beta$. We sometimes use the letter $y_{\beta}$ and $y^{\prime}_{\beta}$ for a highest weight vector and lowest weight vector of $L_{\beta}$, respectively. The Langlands dual $^{L}\mathfrak{g}$ is defined by replacing the root system by the coroot system. Namely,
$(Q_{{}^{L}\mathfrak{g}},(\cdot,\cdot)_{{}^{L}Q})=(Q^{\vee},\tfrac{1}{r^{\vee}}(\cdot,\cdot)_{Q}),$
$\alpha_{i,{}^{L}\mathfrak{g}}=\alpha_{i}^{\vee},\ \ \alpha_{i,{}^{L}\mathfrak{g}}^{\vee}=r^{\vee}\alpha_{i},\ \ \alpha_{i,{}^{L}\mathfrak{g}}^{\ast}=r^{\vee}\varpi_{i},\ \ \varpi_{i,{}^{L}\mathfrak{g}}=\alpha_{i}^{\ast}$
(in particular, $\rho_{{}^{L}\mathfrak{g}}=\rho^{\vee}$ and $\rho_{{}^{L}\mathfrak{g}}^{\vee}=r^{\vee}\rho$). If we want to emphasize that a certain object $X$ (e.g., $\mathfrak{g}$, QQ,…) is derived from $^{L}\mathfrak{g}$ rather than $\mathfrak{g}$, we often use symbols such as $^{L}X$, $_{{}^{L}\mathfrak{g}}, Xˇ\check{X}$, etc. By abuse of notation, for $\theta=\sum_{i\in I}a_{i}\alpha_{i}$, denote $^{L}\theta=\sum_{i\in I}a_{i}\alpha_{r+1-i}^{\vee}\in\mathfrak{h}^{\ast}$ (i.e., we regard $^{L}\theta_{\mathfrak{g}}=\theta_{{}^{L}\mathfrak{g}}$ as an element in $\mathfrak{h}^{\ast}$). For more detailed data in each cases, see e.g. [34, p.91-92] (but note that the labeling of Dynkin diagrams are different).
For a weight $B$-module $M$ and $\beta\in\mathfrak{h}^{\ast}$, denote $M^{h=\beta}$ the weight space with Cartan weight $\beta$. For $\mu\in\mathfrak{h}^{\ast}$, let $\mathbb{C}_{\mu}$ be the one-dimensional $B$-module such that $\mathfrak{n}_{-}$-action is trivial. For a $B$-module $V$ and $\mu\in\mathfrak{h}^{\ast}$, we write $V(\mu)$ for the $B$-module $V\otimes_{\mathbb{C}}\mathbb{C}_{\mu}$. We use the same notation for a vector bundle and its sheaf of sections. For a sheaf $\mathcal{F}$ over a topological space $X$ and an open subset $U$ of $X$, $\mathcal{F}(U)$ denotes the space of sections of $\mathcal{F}(U)$ on $U$. For an algebraic variety $X$, let $\mathcal{O}_{X}$ be the structure sheaf of $X$. For $\mu\in\mathfrak{h}^{\ast}$, we write $\mathcal{O}(\mu)$ for the line bundle (or invertible sheaf) $G\times_{B}\mathbb{C}_{\mu}$ over the flag variety $G/B$. In particular, $\mathcal{O}(0)$ is $\mathcal{O}_{G/B}$. For an $\mathcal{O}_{G/B}$-module $\mathcal{F}$, we use the letter $\mathcal{F}(\mu)$ for $\mathcal{F}\otimes_{\mathcal{O}_{G/B}}\mathcal{O}(\mu)$. In particular, for a $B$-module $V$ and a homogeneous vector bundle $G\times_{B}V$, we have $(G\times_{B}V)(\mu)=G\times_{B}V(\mu)$.
For later convenience, we will list several facts.
Lemma 1.3.
[45, (114), (115), Lemma 4.5] For a $P_{i}$-module $M$ and $\mu\in P$, we have
$H^{n}(P_{i}\times_{B}M(\mu))\simeq H^{n}(P_{i}\times_{B}\mathbb{C}_{\mu})\otimes M,$
$\dim_{\mathbb{C}}H^{n}(P_{i}\times_{B}\mathbb{C}_{\mu})=\begin{cases}(\mu,\alpha_{i}^{\vee})+1&(n=0,\ (\mu,\alpha_{i}^{\vee})\geq 0),\\ -(\mu,\alpha_{i}^{\vee})-1&(n=1,\ (\mu,\alpha_{i}^{\vee})<0),\\ 0&(\text{otherwise}).\end{cases}$
In particular, if $(\mu+\rho,\alpha_{i}^{\vee})\geq 0$, then $H^{0}(P_{i}\times_{B}\mathbb{C}_{\mu})\simeq H^{1}(P_{i}\times_{B}\mathbb{C}_{\sigma_{i}(\mu+\rho)-\rho})$ as $P_{i}$-submodules.
Lemma 1.4.
If $M$ is not only $B$-module but also an $SL_{2}^{j}$-module, then $M$ has a structure of $P_{j}$-module.
Proof.
The isomorphism of varieties $P_{j}/B\simeq SL_{2}^{j}/B^{j}(\simeq\mathbf{P}^{1})$ actually extends to isomorphism of equivariant bundles $P_{j}\times_{B}M\simeq SL_{2}^{j}\times_{B^{j}}M$ over $\mathbf{P}^{1}$ and thus induces an isomorphism of the global sections
[…]
Since the left-hand side implies that it is naturally a $P_{j}$-module, we obtain the assertion. ∎
Lemma 1.5.
Let $J\subset I$ and $M_{j}$ ($j\in J$) be a $P_{j}$-submodule of a $P_{J}$-submodule $M$. If the intersection
$\bigcap_{j\in J}M_{j}\subset M$
is closed under $e_{j}$ ($j\in J$), then it is a $P_{J}$-submodule of $M$.
Proof.
It suffices to show $\mathrm{ad}_{e_{i}}^{1-c_{ij}}(e_{j})=0$ for $i\neq j$ on $\bigcap_{j\in J}M_{j}$. Since $e_{i},f_{j}]=0$, we have
$[\mathrm{ad}_{e_{i}}^{1-c_{ij}}(e_{j}),f_{i}]=\sum_{m=0}^{-c_{ij}}\mathrm{ad}_{e_{i}}^{-c_{ij}-m}\mathrm{ad}_{h_{i}}\mathrm{ad}_{e_{i}}^{m}(e_{j})=\sum_{m=0}^{-c_{ij}}h_{i}(m\alpha_{i}-\alpha_{j})\mathrm{ad}_{e_{i}}^{-c_{ij}}(e_{j})=0$
Then, it suffices to show $\mathrm{ad}_{e_{i}}^{1-c_{ij}}(e_{j})y^{i}_{\alpha}=0$ for any highest weight vector $y^{i}_{\alpha}$ of $\bigcap_{j\in J}M_{j}$ with respect to the $SL_{2}^{i}$-action (where $\alpha\in P$ such that $(\alpha,\alpha_{i}^{\vee})\geq 0$). Since $h_{i}((1-c_{ij})\alpha_{i}+\alpha_{j}+\alpha)=2-c_{ij}+h_{i}(\alpha)$, it suffices to show $f_{i}^{2-c_{ij}+h_{i}(\alpha)}\mathrm{ad}_{e_{i}}^{1-c_{ij}}(e_{j})y^{i}_{\alpha}=0$. Note that $f_{i}^{2-c_{ij}+h_{i}(\alpha)}e_{i}^{1-c_{ij}}=Xf_{i}^{h_{i}(\alpha)+1}$ for some $X\in U(\mathfrak{sl}_{2}^{i})$ by weight consideration. Then
$f_{i}^{2-c_{ij}+h_{i}(\alpha)}\mathrm{ad}_{e_{i}}^{1-c_{ij}}(e_{j})y^{i}_{\alpha}=f_{i}^{2-c_{ij}+h_{i}(\alpha)}e_{i}^{1-c_{ij}}e_{j}y^{i}_{\alpha}=Xf_{i}^{s+1}e_{j}y^{i}_{\alpha}=Xe_{j}f_{i}^{s+1}y^{i}_{\alpha}=0.$
This completes the proof. ∎
1.2. Shift system
In this subsection, we introduce a new concept, named shift system [^6] The name “shift system” was chosen because, as we show below, the shift $\sigma\uparrow\lambda$ plays a central role in our representation theory. However, this is tentative (an alternative name could be FT data, Felder/screening system, etc.) and may be changed to a more appropriate name/concept as our understanding of the theory of FT construction improves., which is the main tool in the present paper, and give its basic examples and properties.
Definition 1.6.
The shift system refers to a triple
$(\Lambda,\uparrow,\{V_{\lambda}\}_{\lambda\in\Lambda})$
that satisfies the following conditions:
  1. (1)
    $\Lambda$ is a $W$-module. Denote $\ast\colon W\times\Lambda\rightarrow\Lambda$ the $W$-action on $\Lambda$.
  2. (2)
    $\uparrow\colon W\times\Lambda\rightarrow P$ is a map (shift map) such that
    1. (a)
      $\sigma_{i}\sigma\uparrow\lambda=\sigma_{i}(\sigma\uparrow\lambda)+\sigma_{i}\uparrow(\sigma\ast\lambda)$.
    2. (b)
      If $l(\sigma_{i}\sigma)=l(\sigma)+1$, then $(\sigma\uparrow\lambda,\alpha_{i}^{\vee})\geq 0$.
    3. (c)
      If $\lambda\not\in\Lambda^{\sigma_{i}}$, then $(\sigma_{i}\uparrow\lambda,\alpha_{i}^{\vee})=-1$. If $\lambda\in\Lambda^{\sigma_{i}}$, then $\sigma_{i}\uparrow\lambda=-\alpha_{i}$.
  3. (3)
    $V_{\lambda}$ is a weight $B$-module such that
    1. (a)
      $V_{\lambda}=\bigoplus_{\Delta}V_{\lambda,\Delta}$ and each $V_{\lambda,\Delta}$ is a finite-dimensional weight $B$-submodule.
    2. (b)
      For any $i\in I$ and $\lambda\not\in\Lambda^{\sigma_{i}}$, there exists $P_{i}$-submodules $W_{i,\lambda}\subseteq V_{\lambda}$, $W_{i,\sigma_{i}\ast\lambda}\subseteq V_{\sigma_{i}\ast\lambda}$ and a $B$-module homomorphism $Q_{i,\lambda}\colon V_{\lambda}\rightarrow W_{i,\sigma_{i}\ast\lambda}({\sigma_{i}}\uparrow{\lambda})$ such that we have the short exact sequence
      $0\rightarrow W_{i,\lambda}\rightarrow V_{\lambda}\xrightarrow{Q_{i,\lambda}}W_{i,\sigma_{i}\ast\lambda}(\sigma_{i}\uparrow\lambda)\rightarrow 0$
      of $B$-modules. If $\lambda\in\Lambda^{\sigma_{i}}$, then $V_{\lambda}$ has the $P_{i}$-module structure (we sometimes write $W_{i,\lambda}=V_{\lambda}$).
Let us introduce some terms for convenience. The action of $f_{i}\in\mathfrak{n}_{-}$ on $V_{\lambda}$ and the $B$-module homomorphism $Q_{i,\lambda}$ above are referred to as a long screening operator and a short screening operator, respectively. If some object $X$ is isomorphic (resp. conjectured to be isomorphic) to $H^{0}(P\times_{B}V_{\lambda})$ for some $V_{\lambda}$ and a parabolic subgroup $P\supseteq B$, we call $H^{0}(P\times_{B}V_{\lambda})$ the Feigin–Tipunin construction of $X$ (resp. FT conjecture on $X$).
Remark 1.7.
Let us list several easy facts on the shift system to check.
  1. (1)
    By substituting $\sigma=\operatorname{id}$ to the axiom (2a), we have $\operatorname{id}\uparrow\lambda=0$. On the other hand, substituting $\sigma=\sigma_{i}$ shows that $\sigma_{i}\uparrow\lambda+\sigma_{i}\uparrow(\sigma_{i}\ast\lambda)=-\alpha_{i}$ ($\lambda\not\in\Lambda^{\sigma_{i}}$) or $2\alpha_{i}$ ($\lambda\in\Lambda^{\sigma_{i}}$). By axioms (2a) and (2b), if $l(\sigma_{i}\sigma)=l(\sigma)-1$, then $(\sigma\uparrow\lambda,\alpha_{i}^{\vee})\in\mathbb{Z}_{<0}$.
  2. (2)
    For $\sigma\in W$, let us fix a minimal expression $\sigma=\sigma_{i_{m}}\cdots\sigma_{i_{1}}$. Repeated use of axiom (2a) shows that
    $\sigma\uparrow\lambda=(\sigma_{i_{m}}\cdots\sigma_{i_{1}})\uparrow\lambda=\sum_{j=1}^{m}\sigma_{i_{j}}\uparrow(\sigma_{i_{j-1}}\cdots\sigma_{i_{0}}\ast\lambda)-\sum_{j=1}^{m}(\sigma_{i_{j-1}}\cdots\sigma_{i_{1}}\uparrow\lambda,\alpha_{i_{j}}^{\vee})\alpha_{i_{j}}.$
    By combining it with axiom (2b), the condition (strong) is equivalent to the following:
    (1.3) For any $1\leq m\leq n$, $(\sigma_{i_{m}}\cdots\sigma_{i_{1}})\uparrow\lambda=\sum_{j=1}^{m}\sigma_{i_{j}}\uparrow(\sigma_{i_{j-1}}\cdots\sigma_{i_{0}}\ast\lambda)$.
    Note that at this point, these conditions depend on the minimal expression of $\sigma$. In Example 1.8 below, the independence of the choice of minimal expression is clear. Note that if $w_{0}\uparrow\lambda$ is independent of the minimal expression of $w_{0}$, then we have $w_{0}\uparrow\lambda=-\rho$.
  3. (3)
    The $P_{i}$-submodule $W_{i,\lambda}$ above is the maximal $P_{i}$-submodule of $V_{\lambda}$. In fact, by applying the long exact sequence $H^{\bullet}(P_{i}\times_{B}-)$ to the short exact sequence, we have $H^{0}(P_{i}\times_{B}V_{\lambda})\simeq W_{i,\lambda}$. If $M$ is a $P_{i}$-submodule of $V_{\lambda}$, then by Lemma 1.3, we have the $P_{i}$-module isomorphism
    $M\simeq H^{0}(P_{i}\times_{B}M)\subseteq H^{0}(P_{i}\times_{B}V_{\lambda})\simeq W_{i,\lambda},$
    where both isomorphisms above are given by the evaluation map $\operatorname{ev}\colon s\mapsto s(\operatorname{id}_{P_{i}/B})$ and its inverse.
  4. (4)
    For $j\in I$, $\beta\in P$ such that $(\beta,\alpha_{j}^{\vee})\geq 0$ and $v\in V_{\lambda}^{h=\beta}$, we have $v=0$ (resp. $v\in W_{j,\lambda}$) iff $f_{j}^{(\beta,\alpha_{j}^{\vee})}v=0$ (resp. $f_{j}^{(\beta,\alpha_{j}^{\vee})+1}v=0$).
Example 1.8.
Let us give an example of $(\Lambda,\uparrow)$ in the shift system. For a fixed $x\in\mathfrak{h}^{\ast}_{\mathbb{R}}$, any $\mu\in\mathfrak{h}^{\ast}_{\mathbb{R}}$ has the unique decomposition $\mu=-\mu^{\bullet}+\mu_{\bullet}$, where $\mu^{\bullet}\in P$ and $\mu_{\bullet}\in\mathfrak{h}^{\ast}_{\mathbb{R}}$ such that $0<(\mu_{\bullet}+x,\alpha_{i}^{\vee})\leq 1$ for any $i\in I$. For $\sigma\in W$, set $\sigma\ast\mu=-\mu^{\bullet}+\sigma(\mu_{\bullet}+x)-x$. It defines a $W$-action on $\mathfrak{h}^{\ast}/Q$, and let $\Lambda$ be the unique representatives of a $W$-submodule of $\mathfrak{h}^{\ast}/Q$ (i.e., in the unique decomposition $\lambda=-\lambda^{\bullet}+\lambda_{\bullet}\in\Lambda$, $\lambda^{\bullet}\in P_{\mathrm{min}}$). Then the $\Lambda$ and the “carry-over of the $W$-action”
$\sigma\uparrow\lambda:=\sigma\ast\lambda_{\bullet}-(\sigma\ast\lambda)_{\bullet}\in P$
satisfies the axiom in Definition 1.6 (well-definedness and the independence above are clear). In fact, we have
$\sigma_{i}\sigma\uparrow\lambda$ $=\sigma_{i}\sigma\ast\lambda_{\bullet}-({\sigma_{i}\sigma\ast\lambda})_{\bullet}$
$=\sigma_{i}\sigma(\lambda_{\bullet}+x)-x-({\sigma_{i}\sigma\ast\lambda})_{\bullet}$
$=\sigma(\lambda_{\bullet}+x)-x-(\sigma(\lambda_{\bullet}+x),\alpha_{i}^{\vee})\alpha_{i}-({\sigma_{i}\sigma\ast\lambda})_{\bullet}$
$=\sigma\ast\lambda_{\bullet}-({\sigma\ast\lambda})_{\bullet}+({\sigma\ast\lambda})_{\bullet}-(\sigma(\lambda_{\bullet}+x)-x-({\sigma\ast\lambda})_{\bullet}+({\sigma\ast\lambda})_{\bullet}+x,\alpha_{i}^{\vee})\alpha_{i}-({\sigma_{i}\sigma\ast\lambda})_{\bullet}$
$=\sigma\uparrow\lambda+({\sigma\ast\lambda})_{\bullet}-(\sigma\uparrow\lambda+({\sigma\ast\lambda})_{\bullet}+x,\alpha_{i}^{\vee})\alpha_{i}-({\sigma_{i}\sigma\ast\lambda})_{\bullet}$
$=\sigma_{i}(\sigma\uparrow\lambda)+\sigma_{i}\uparrow(\sigma\ast\lambda),$
and thus (2a) is satisfied. Furthermore, if $l(\sigma_{i}\sigma)=l(\sigma)+1$, then we have
$(\sigma_{i}\sigma\uparrow\lambda,\alpha_{i}^{\vee})=(\sigma_{i}\sigma\ast\lambda_{\bullet}-({\sigma_{i}\sigma\ast\lambda})_{\bullet},\alpha_{i}^{\vee})=-(\sigma(\lambda_{\bullet}+x),\alpha_{i}^{\vee})-(({\sigma_{i}\sigma\ast\lambda})_{\bullet}+x,\alpha_{i}^{\vee})\in\mathbb{Z}_{<0},$
because $\sigma^{-1}\alpha_{i}\in\Delta_{+}$. In particular, when $\sigma=\operatorname{id}$, we have
$(\sigma_{i}\uparrow\lambda,\alpha_{i}^{\vee})=-(\lambda_{\bullet}+x,\alpha_{i}^{\vee})-(({\sigma_{i}\ast\lambda})_{\bullet}+x,\alpha_{i}^{\vee})\in\{-1,-2\}$
and thus (2c) follows from the assumption. On the other hand, since
$(\sigma_{i}\sigma\uparrow\lambda,\alpha_{i}^{\vee})=(\sigma_{i}(\sigma\uparrow\lambda)+\sigma_{i}\uparrow(\sigma\ast\lambda),\alpha_{i}^{\vee})=-(\sigma\uparrow\lambda,\alpha_{i}^{\vee})+(\sigma_{i}\uparrow(\sigma\ast\lambda),\alpha_{i}^{\vee}),$
we have $(\sigma\uparrow\lambda,\alpha_{i}^{\vee})>(\sigma_{i}\uparrow(\sigma\ast\lambda),\alpha_{i}^{\vee})$. By the assumption and the discussion above, we can check (2b)
For a shift system of this example, we use the notation
(1.4) $H^{0}(G\times_{B}V_{\lambda})\simeq\bigoplus_{\alpha\in P_{+}\cap Q}L_{\alpha+\lambda^{\bullet}}\otimes\mathcal{W}_{-\alpha+\lambda}.$
In Part II, we will consider the cases where $\Lambda$ is the unique representative of $\tfrac{1}{p}Q^{\ast}/Q$ for the $W$-action defined by $x=\tfrac{1}{p}\rho^{\vee}$ ($p=r^{\vee}m$, $m\in\mathbb{Z}_{\geq 1}$) and $x=\tfrac{1}{p}\rho$ (p=2m1p=2m-1, $m\in\mathbb{Z}_{\geq 1}$ and $\mathfrak{g}=B_{r}$), respectively.
The following lemma is shown in exactly the same manner as [45], so the proof is omitted.
Lemma 1.9.
[45, Lemma 3.17-3.18] Let $x\in V_{\lambda}^{h=\beta}$ be a nonzero vector for some $\beta\in P_{+}$. If there exists a $B$-module homomorphism $\Phi\colon L_{\beta}\rightarrow U(\mathfrak{b})x$ that sends the highest weight vector $y_{\beta}$ of $L_{\beta}$ to $x$, then $\Phi$ is an isomorphism. In particular, $U(\mathfrak{b})x$ is a $G$-submodule of $V_{\lambda}$.
1.3. Feigin–Tipunin conjecture/construction
Let $(\Lambda,\uparrow,\{V_{\lambda}\}_{\lambda\in\Lambda})$ be a shift system. In this subsection, we will give a brief necessary and sufficient condition for the Feigin-Tipunin conjecture [27, 45]
$H^{0}(G\times_{B}V_{\lambda})\simeq\bigcap_{i\in I}W_{i,\lambda}\subseteq V_{\lambda}$
to hold. For $(i,j,\lambda)\in I\times I\times\Lambda$, we consider the following condition (see [45, (98),(99),(100)]):
(1.5) $\text{$\lambda\in\Lambda^{\sigma_{j}}$ or $({\sigma_{j}}\uparrow{\lambda},\alpha_{i}^{\vee})=-\delta_{ij}$}.$
For a subset $J\subseteq I$, we also consider the following condition:
[…]
When $J=I$, the condition (1.6) is stated as follows (this is the weak condition (weak) in Theorem 1.1(1)):
(1.7) $(i,j,\lambda)$ satisfies (1.5) for any $(i,j)\in I\times I$.
Lemma 1.10.
[45, Lemma 4.3] Let $i,j\in I$ and $\lambda\in\Lambda$. If $(i,j,\lambda)$ satisfies (1.5), then $W_{i,\lambda}\cap W_{j,\lambda}$ is closed under the operator $e_{i}$. In particular, if $(J,\lambda)\subseteq I\times\Lambda$ satisfies (1.6), then $\bigcap_{j\in J}W_{j,\lambda}$ is a $P_{J}$-module.
Proof.
If $\lambda\in\Lambda^{\sigma_{j}}$ or $i=j$, then we have $W_{i,\lambda}\cap W_{j,\lambda}=W_{i,\lambda}$ and the assertion is clear. Let us assume that $i\not=j$ and $({\sigma_{j}}\uparrow{\lambda},\alpha_{i}^{\vee})=0$. It suffices to show that if $f_{i}^{n}A_{\beta}\in W_{i,\lambda}\cap W_{j,\lambda}$ for some $A_{\beta}\in(W_{i,\lambda})_{\beta}^{e_{i}}$, $(\beta,\alpha_{i}^{\vee})>0$, and $0<n\leq(\beta,\alpha_{i}^{\vee})$, then $A_{\beta}$ (and thus $f_{i}^{n-1}A_{\beta}$) is also in $W_{i,\lambda}\cap W_{j,\lambda}$. Since $f_{i}^{(\beta,\alpha_{i}^{\vee})}A_{\beta}\in W_{j,\lambda}$, we have
(1.8) $f_{i}^{(\beta,\alpha_{i}^{\vee})}Q_{j,\lambda}A_{\beta}=Q_{j,\lambda}f_{i}^{(\beta,\alpha_{i}^{\vee})}A_{\beta}=0,$
By the assumption $({\sigma_{j}}\uparrow{\lambda},\alpha_{i}^{\vee})=0$ and Remark 1.7(4), we have $Q_{j,\lambda}A_{\beta}=0$, that is, $A_{\beta}\in W_{j,\lambda}$. The last assertion follows from Lemma 1.4 and Lemma 1.5. ∎
Lemma 1.11.
[45, Theorem 4.4] Only if $(J,\lambda)$ satisfies (1.6), then $\bigcap_{j\in J}W_{j,\lambda}$ is a $P_{J}$-submodule of $V_{\lambda}$.
Proof.
Let us assume that $(J,\lambda)$ does not satisfy (1.6). Then there exists a pair $(j,i)\in J\times J$ such that $i\not=j$, $\sigma_{i}\uparrow\lambda\not=-\alpha_{i}$ and $(\sigma_{i}\uparrow\lambda,\alpha_{j}^{\vee})>0$. For $\beta\in P$ such that $(\beta,\alpha_{k}^{\vee})\geq 0$ for $k\in J$, we take nonzero vectors
$x\in\bigcap_{k\in J}(W_{k,\sigma\ast\lambda}^{h=\beta})^{e_{k}},\ \ y=f_{i}^{(\beta,\alpha_{i}^{\vee})}x,\ \ z=f_{j}^{\sigma_{i}(\beta),\alpha_{j}}y.$
Clearly we have $f_{i}y=0$. Furthermore, since $y\in(W_{j,\sigma_{i}\ast\lambda}^{h=\sigma_{i}(\beta)})^{e_{j}}$, we have $f_{j}z=0$. Denote $w\in V_{\lambda}^{h=\beta+\sigma_{i}\uparrow\lambda}$ is a preimage of $x$, namely, we have $Q_{i,\lambda}w=x$. Then
(1.9) $0\not=f_{i}^{(\beta,\alpha_{i}^{\vee})}w\in\bigcap_{i\not=k\in J}(W_{k,\lambda}^{h=\sigma(\beta)+\sigma_{i}\uparrow\lambda})^{e_{k}}.$
In fact, using the Serre relation $\mathrm{ad}(f_{k})^{1-(\alpha_{k}^{\vee},\alpha_{i})}f_{i}=0$ repeatedly, we have
$f_{k}^{(\sigma_{i}(\beta)+\sigma_{i}\uparrow\lambda,\alpha_{k}^{\vee})+1}f_{i}^{(\beta,\alpha_{i}^{\vee})}w=f_{k}^{(\beta+\sigma_{i}\uparrow\lambda,\alpha_{k}^{\vee})+1}(f_{k}^{-(\alpha_{k}^{\vee},\alpha_{i})(\beta,\alpha_{i}^{\vee})}f_{i}^{(\beta,\alpha_{i}^{\vee})})w=0.$
Let us define the nonzero vector $z^{\prime}$ by
$z^{\prime}=f_{j}^{(\sigma_{i}(\beta)+\sigma_{i}\uparrow\lambda,\alpha_{j}^{\vee})}f_{i}^{(\beta,\alpha_{i}^{\vee})}w.$
Then by the assumption $(\sigma_{i}\uparrow\lambda,\alpha_{j}^{\vee})>0$, we have
$Q_{i,\lambda}z^{\prime}=f_{j}^{(\sigma_{i}(\beta)+\sigma_{i}\uparrow\lambda,\alpha_{j}^{\vee})}Q_{i,\lambda}f_{i}^{(\beta,\alpha_{i}^{\vee})}w=f_{j}^{(\sigma_{i}(\beta)+\sigma_{i}\uparrow\lambda,\alpha_{j}^{\vee})}y=f_{j}^{(\sigma_{i}\uparrow\lambda,\alpha_{j}^{\vee})}z=0.$
By (1.9), we have $z^{\prime}\in\bigcap_{k\in J}W_{k,\lambda}$. On the other hand, if $\bigcap_{k\in J}W_{k,\lambda}$ is closed under $e_{j}$, then we have
$f_{i}^{(\beta,\alpha_{i}^{\vee})}w\in\bigcap_{k\in J}W_{k,\lambda},$
and thus the image $y$ of $Q_{i,\lambda}$ is zero. It contradicts to the fact $y\not=0$ above, and thus $\bigcap_{k\in J}W_{k,\lambda}$ is not closed under $e_{j}$, and hence not a $P_{J}$-submodule of $V_{\lambda}$. ∎
Theorem 1.12.
[45, Theorem 4.14] The evaluation map
$\mathrm{ev}\colon H^{0}(G\times_{B}V_{\lambda})\rightarrow V_{\lambda},\ \ s\mapsto s(\operatorname{id}_{G/B})$
is an injective $B$-module homomorphism. In particular, $H^{0}(G\times_{B}V_{\lambda})$ sends to the maximal $G$-submodule of $V_{\lambda}$.
Proof.
Since the proof is the same as [45, Lemma 4.15 - 4.18], we omit the detail. For $\beta\in P_{+}$, denote $W_{\beta}$ the set of highest weight vectors of $H^{0}(G\times_{B}V_{\lambda})$ with highest weight $\beta$. Let $G/B=\bigcap_{\sigma\in W}U_{\sigma}$ be the Schubert open covering. For $s\in W_{\beta}$, we have $s(N_{+})=s(\operatorname{id})$, and thus $s|_{U_{\operatorname{id}}}=1\otimes v$ for some $v\in V_{\lambda}^{h=\beta}$. By considering coordinate changes between $U_{\sigma}$’s, $\operatorname{ev}|_{W_{\beta}}$ is injective. By Lemma 1.9, $\operatorname{ev}|_{L_{\beta}\otimes W_{\beta}}$ so is. Now we just need to make sure that $L_{\beta}\otimes W_{\beta}$’s are linearly independent with respect to different highest weights $\beta$’s. For a fixed $\gamma\in P$ and $\Delta$, let us take a vector $s=\sum_{\beta\in P_{+}}s_{\beta}\in H^{0}(G\times_{B}V_{\lambda,\Delta})^{h=\gamma}$ such that $\operatorname{ev}(s)=0$. We fix a minimal highest weight $\alpha\in P_{+}$ in $\{\beta\in P_{+}\ |\ s_{\beta}\not=0\}$ with respect to the standard partial order $\geq$ on $\mathfrak{h}^{\ast}$, i.e. $\mu\geq\mu^{\prime}$ iff $\mu-\mu^{\prime}\in\bigoplus_{i\in I}\mathbb{Z}_{\geq 0}\alpha_{i}$. Since $L_{\beta_{2}}^{h=w_{0}(\beta_{1})}=\{0\}$ for $\beta_{1}\not\leq\beta_{2}$, there exists $X\in U(\mathfrak{b})$ such that
$0=X\operatorname{ev}(s)=X\operatorname{ev}(s_{\alpha})=\operatorname{ev}(Xs_{\alpha}).$
As $\operatorname{s}|_{L_{\alpha}\otimes W_{\alpha}}$ is injective, we have $s_{\alpha}=0$. By repeating this procedure, we have $s=0$ and thus $\operatorname{ev}$ is injective. ∎
Let us prove Theorem 1.1(1) (see also [45, Section 4.5]). By Theorem 1.12 and Remark 1.7(3), the first half of Theorem 1.1(1) is proved. By Lemma 1.10 and Lemma 1.11, $\bigcap_{i\in I}W_{i,\lambda}$ has the $G$-module structure (and thus, in $H^{0}(G\times_{B}V_{\lambda})$) if and only if $\lambda\in\Lambda$ satisfies the weak condition (weak). This completes the proof.
1.4. Borel–Weil–Bott-type theorem
Let $(\Lambda,\uparrow,\{V\}_{\lambda\in\Lambda})$ be a shift system. For $\mu\in P$ and $\lambda\not\in\Lambda^{\sigma_{i}}$, we have
(1.10) $0\rightarrow W_{i,\lambda}(\mu)\rightarrow V_{\lambda}(\mu)\rightarrow W_{i,\sigma_{i}\ast\lambda}(\mu+\sigma_{i}\uparrow\lambda)\rightarrow 0.$
Lemma 1.13.
[45, Lemma 4.10] For $i\in I$, $\sigma\in W$, $\lambda\in\Lambda$ such that $\ell(\sigma\sigma_{i})=\ell(\sigma)+1$ and $\sigma\ast\lambda\not\in\Lambda^{\sigma_{i}}$, we have short exact sequences of $P_{i}$-modules
$0\rightarrow H^{0}(P_{i}\times_{B}\mathbb{C}_{\sigma\uparrow\lambda})\otimes W_{i,\sigma\ast\lambda}\rightarrow H^{0}(P_{i}\times_{B}V_{\sigma\ast\lambda}(\sigma\uparrow\lambda))\rightarrow H^{0}(P_{i}\times_{B}\mathbb{C}_{\sigma\uparrow\lambda+\sigma_{i}\uparrow(\sigma\ast\lambda)})\otimes W_{i,\sigma_{i}\sigma\ast\lambda}\rightarrow 0,$
$0$ $\rightarrow H^{1}(P_{i}\times_{B}\mathbb{C}_{\sigma_{i}\sigma\uparrow\lambda})\otimes W_{i,\sigma_{i}\sigma\ast\lambda}$
$\rightarrow H^{1}(P_{i}\times_{B}V_{\sigma_{i}\sigma\ast\lambda}(\sigma_{i}\sigma\uparrow\lambda))$
$\rightarrow H^{1}(P_{i}\times_{B}\mathbb{C}_{\sigma_{i}\sigma\uparrow\lambda+\sigma_{i}\uparrow(\sigma_{i}\sigma\ast\lambda)})\otimes W_{i,\sigma\ast\lambda}\rightarrow 0.$
Proof.
Let us apply the long exact sequence $H^{\bullet}(P_{i}\times_{B}\text{-})$ for (1.10) with
$(\lambda,\mu)\mapsto(\sigma\ast\lambda,\sigma\uparrow\lambda),\quad(\lambda,\mu)\mapsto(\sigma_{i}\sigma\ast\lambda,(\sigma_{i}\sigma)\uparrow\lambda)$
respectively. Then the assertions follow from Lemma 1.3 if we have
$H^{1}(P_{i}\times_{B}\mathbb{C}_{\sigma\uparrow\lambda})=H^{0}(P_{i}\times_{B}\mathbb{C}_{\sigma_{i}\sigma\uparrow\lambda+\sigma_{i}\uparrow(\sigma_{i}\sigma\ast\lambda)})=0.$
By Lemma 1.3 again, it suffices to show that
$(\sigma\uparrow\lambda,\alpha_{i}^{\vee})\geq 0,\ (\sigma_{i}\sigma\uparrow\lambda+\sigma_{i}\uparrow(\sigma_{i}\sigma\ast\lambda),\alpha_{i}^{\vee})<0.$
They immediately follow from the assumption and axioms in Definition 1.6. ∎
Theorem 1.14.
[45, Theorem 4.8] Let us assume that $V_{\lambda}$ is a $P_{i}$-module if $\lambda\in\Lambda^{\sigma_{i}}$. For $\lambda\in\Lambda$ and $\sigma\in W$ satisfying $\ell(\sigma_{i}\sigma)=\ell(\sigma)+1$ and $(\sigma\uparrow\lambda,\alpha_{i}^{\vee})=0$, we have an isomorphism
(1.11) $H^{n}(G\times_{B}V_{\sigma\ast\lambda}(\sigma\uparrow\lambda))\simeq H^{n+1}(G\times_{B}V_{\sigma_{i}\sigma\ast\lambda}(\sigma_{i}\sigma\uparrow\lambda)).$
In particular, Theorem 1.1(2) is proved by applying this isomorphism repeatedly.
Proof.
First we consider the case $\sigma\ast\lambda\not\in\Lambda^{\sigma_{i}}$. By the axiom (2c) and the assumption $(\sigma\uparrow\lambda,\alpha_{i}^{\vee})=0$, we have $(\sigma\uparrow\lambda+\sigma_{i}\uparrow(\sigma\ast\lambda),\alpha_{i}^{\vee})=-1$. Therefore, by Lemma 1.3, we have
(1.12) $H^{n}(P_{i}\times_{B}\mathbb{C}_{\sigma\uparrow\lambda})\simeq\delta_{n,0}\mathbb{C}_{\sigma\uparrow\lambda},\quad H^{n}(P_{i}\times_{B}\mathbb{C}_{\sigma\uparrow\lambda+\sigma_{i}\uparrow(\sigma\ast\lambda)})=0$
and by applying (1.12) to the first short exact sequence in Lemma 1.13, we have
(1.13) $H^{n}(P_{i}\times_{B}V_{\sigma\ast\lambda}(\sigma\uparrow\lambda))\simeq\delta_{n,0}W_{i,\sigma\ast\lambda}(\sigma\uparrow\lambda).$
By the assumption $\sigma\ast\lambda\not\in\Lambda^{\sigma_{i}}$ (and thus, $\sigma_{i}\sigma\ast\lambda\not\in\Lambda^{\sigma_{i}}$) and axiom (2c), we have
(1.14) $\sigma_{i}\sigma\uparrow\lambda=\sigma_{i}\circ(\sigma\uparrow\lambda+\sigma_{i}\uparrow(\sigma\ast\lambda)),~\sigma\uparrow\lambda=\sigma_{i}\circ(\sigma_{i}\sigma\uparrow\lambda+\sigma_{i}\uparrow(\sigma_{i}\sigma\ast\lambda)).$
By applying Lemma 1.3 to (1.14), we obtain that
(1.15) $H^{n}(P_{i}\times_{B}\mathbb{C}_{\sigma_{i}\sigma\uparrow\lambda})=0,~H^{n}(P_{i}\times_{B}\mathbb{C}_{\sigma_{i}\sigma\uparrow\lambda+\sigma_{i}\uparrow(\sigma_{i}\sigma\ast\lambda)})\simeq\delta_{n,1}\mathbb{C}_{\sigma\uparrow\lambda},$
where the first one follows from the assumption $(\sigma\uparrow\lambda,\alpha_{i}^{\vee})=0$ and (2c), and the second one follows from $(\sigma\uparrow\lambda,\alpha_{i}^{\vee})=0$ and Lemma 1.3. Hence, by combining the second short exact sequence with (1.15), we obtain
(1.16) $H^{n}(P_{i}\times_{B}V_{\sigma_{i}\sigma\ast\lambda}(\sigma_{i}\sigma\uparrow\lambda))\simeq\delta_{n,1}W_{i,\sigma\ast\lambda}(\sigma\uparrow\lambda).$
By combining (1.13) and (1.16), we obtain
(1.17) $H^{a}(P_{i}\times_{B}V_{\sigma\ast\lambda}(\sigma\uparrow\lambda))\simeq\delta_{a,0}W_{i,\sigma\ast\lambda}(\sigma\uparrow\lambda)\simeq H^{b}(P_{i}\times_{B}V_{\sigma_{i}\sigma\ast\lambda}(\sigma_{i}\sigma\uparrow\lambda))$
for $(a,b)=(0,1)$, $(1,0)$. Now, we have the Leray spectral sequences
$E_{2}^{a,b}=H^{a}(G\times_{P_{i}}H^{b}(P_{i}\times_{B}V_{\sigma\ast\lambda}(\sigma\uparrow\lambda)))\Rightarrow H^{a+b}(G\times_{B}V_{\sigma\ast\lambda}(\sigma\uparrow\lambda)),$
$E_{2}^{a,b}=H^{a}(G\times_{P_{i}}H^{b}(P_{i}\times_{B}V_{\sigma_{i}\sigma\ast\lambda}(\sigma_{i}\sigma\uparrow\lambda)))\Rightarrow H^{a+b}(G\times_{B}V_{\sigma_{i}\sigma\ast\lambda}(\sigma_{i}\sigma\uparrow\lambda)),$
which differ by the 1-shift for $b$ by (1.17). Thus, we obtain the assertion.
Second, we consider the case where $\sigma\ast\lambda\in\Lambda^{\sigma_{i}}$. By the axiom (2c), we have $\sigma_{i}\sigma\uparrow\lambda=\sigma_{i}\circ(\sigma\uparrow\lambda)$. By combining Lemma 1.3 with the assumption that $V_{\sigma\ast\lambda}$ (and thus, $V_{\sigma_{i}\sigma\ast\lambda}$) is a $P_{i}$-module, we have
(1.18) $\begin{split}H^{a}(P_{i}\times_{B}V_{\sigma\ast\lambda}(\sigma\uparrow\lambda))&\simeq H^{a}(P_{i}\times_{B}\mathbb{C}_{\sigma\uparrow\lambda})\otimes V_{\sigma\ast\lambda}\\ &\simeq H^{b}(P_{i}\times_{B}\mathbb{C}_{\sigma_{i}\sigma\uparrow\lambda})\otimes V_{\sigma_{i}\sigma\ast\lambda}\\ &\simeq H^{b}(P_{i}\times_{B}V_{\sigma_{i}\sigma\ast\lambda}(\sigma_{i}\sigma\uparrow\lambda))\end{split}$
for $(a,b)=(0,1)$, $(1,0)$. By applying the assumption $(\sigma\uparrow\lambda,\alpha_{i}^{\vee})=0$ and Lemma 1.3 to (1.18), we have
(1.19) $H^{1}(P_{i}\times_{B}V_{\sigma\ast\lambda}(\sigma\uparrow\lambda))\simeq H^{0}(P_{i}\times_{B}V_{\sigma_{i}\sigma\ast\lambda}(\sigma_{i}\sigma\uparrow\lambda))\simeq 0.$
By combining the Leray spectral sequences above with (1.18) and (1.19), we obtain the assertion. ∎
1.5. Weyl-type character formula
For a graded vector space $V=\bigoplus_{\Delta}V_{\Delta}$, $\dim_{\mathbb{C}}V_{\Delta}<\infty$, denote
$\operatorname{ch}_{q}V=\sum_{\Delta}\dim_{\mathbb{C}}V_{\Delta}q^{\Delta}$
the character (where $q$ and $\Delta$ can take multiple variables, i.e. $q=(q_{i})_{1\leq i\leq n}$, $\Delta=(\Delta_{i})_{1\leq i\leq n}$, $q^{\Delta}=\Pi_{i=1}^{n}q^{\Delta_{i}}$). In the same manner as [45, Section 4.4], by combining the cohomology vanishing777 Note that the Kempf’s vanishing theorem [32, II,4] is proved under a weaker condition than BWB theorem [32, II,5]. The same situation might hold in our case, but the verification of this is a future work (see also the footnote in Section 2.2). $H^{n>0}(G\times_{B}V_{\lambda})=0$ and the Atiyah–Bott localization formula [3], we obtain the Weyl-type character formula of the FT construction $H^{0}(G\times_{B}V_{\lambda})$. An application of the case $n=0$ in Theorem 1.1(2) will be discussed in Section 2.2.
Corollary 1.15.
For $\lambda\in\Lambda$ satisfying the condition (strong), we have
$\operatorname{ch}_{q}H^{0}(G\times_{B}V_{\lambda})=\sum_{\beta\in P_{+}}\dim L_{\beta}\sum_{\sigma\in W}(-1)^{l(\sigma)}\operatorname{ch}_{q}V_{\lambda}^{h=\sigma\circ\beta}=\sum_{\beta\in P_{+}}\dim L_{\beta}\sum_{\sigma\in W}(-1)^{l(\sigma)}\operatorname{ch}_{q}V_{\sigma\ast\lambda}^{h=\beta-\sigma\uparrow\lambda}$
In particular, under the decomposition (1.4), we have
(1.20) $\operatorname{ch}\mathcal{W}_{-\alpha+\lambda}=\sum_{\sigma\in W}(-1)^{l(\sigma)}\operatorname{ch}_{q}V_{\lambda}^{h=\sigma\circ(\alpha+\lambda^{\bullet})}=\sum_{\sigma\in W}(-1)^{l(\sigma)}\operatorname{ch}_{q}V_{\sigma\ast\lambda}^{h=\alpha+\lambda^{\bullet}-\sigma\uparrow\lambda}$
Proof.
As $l(\sigma)=l(\sigma^{-1})$, it suffices to check that $\operatorname{ch}_{q}V_{\lambda}^{h=\sigma\circ\beta}=\operatorname{ch}_{q}V_{\sigma^{-1}\ast\lambda}^{h=\beta-\sigma^{-1}\uparrow\lambda}$ for $\beta\in P$. First let us consider the case $\sigma=\sigma_{i}$. By applying the short exact sequence in Definition 1.6(3), we have
$\operatorname{ch}_{q}V_{\lambda}^{h=\sigma_{i}\circ\beta}=$ $\operatorname{ch}_{q}W_{i,\lambda}^{h=\sigma_{i}\circ\beta}+\operatorname{ch}_{q}W_{i,\sigma_{i}\ast\lambda}^{h=\sigma_{i}\circ\beta-\sigma_{i}\uparrow\lambda}$
$=$ $\operatorname{ch}_{q}W_{i,\lambda}^{h=\sigma_{i}(\sigma_{i}\circ\beta)}+\operatorname{ch}_{q}W_{i,\sigma_{i}\ast\lambda}^{h=\sigma_{i}(\sigma_{i}\circ\beta-\sigma_{i}\uparrow\lambda)}$
$=$ $\operatorname{ch}_{q}W_{i,\sigma_{i}\ast\lambda}^{h=\beta-\sigma_{i}\uparrow\lambda}+\operatorname{ch}_{q}W_{i,\lambda}^{h=\beta-\alpha_{i}=(\beta-\sigma_{i}\uparrow\lambda)-\sigma_{i}\uparrow(\sigma_{i}\ast\lambda)}=\operatorname{ch}_{q}V_{\sigma_{i}\ast\lambda}^{h=\beta-\sigma_{i}\uparrow\lambda},$
where the second equality follows from the fact that $W_{i,\lambda}$ and $W_{i,\sigma_{i}\ast\lambda}$ are $P_{i}$-modules, and the third one follows from the axiom Definition 1.6(2) and Remark 1.7(1). Let us take a minimal expression $\sigma=\sigma_{i_{n}}\cdots\sigma_{i_{1}}$. By applying the same relations with respect to $\sigma_{i_{1}},\sigma_{i_{2}},\dots$ repeatedly, we have
$\operatorname{ch}_{q}V_{\lambda}^{h=\sigma\circ\beta=\sigma_{i_{n}}\circ(\sigma_{i_{n-1}}\cdots\sigma_{i_{1}}\circ\beta)}=$ $\operatorname{ch}_{q}V_{\sigma_{i_{n}}\ast\lambda}^{h=\sigma_{i_{n-1}}\cdots\sigma_{i_{1}}\circ\beta-\sigma_{i_{n}}\uparrow\lambda}$
$=$ $\operatorname{ch}_{q}V_{\sigma_{i_{n-1}}\sigma_{i_{n}}\ast\lambda}^{h=\sigma_{i_{n-2}}\cdots\sigma_{i_{1}}\circ\beta-\sigma_{i_{n}}\uparrow\lambda-\sigma_{i_{n-1}}\uparrow(\sigma_{i_{n}}\ast\lambda)}$
$=$ $\cdots$
$=$ $\operatorname{ch}_{q}V_{\sigma^{-1}\ast\lambda}^{h=\beta-\sum_{j=1}^{n}\sigma_{n+1-j}\uparrow(\sigma_{n+2-j}\cdots\sigma_{i_{n+1}}\ast\lambda)}$
where $\sigma_{i_{n+1}}:=\operatorname{id}$. By Remark 1.7(2), the power of the most right hand side is $\beta-\sigma^{-1}\uparrow\lambda$. ∎
Part II Application to vertex operator superalgebras
In Part II, we consider several free-field algebras and check the axioms of shift system (Definition 1.6). It enables us to use the results in Part I for the study of the corresponding new multiplet W-(super)algebras.
2. Shift systems and free field algebras
2.1. Preliminary from VOSA
First, let us introduce the vertex operator superalgebra and some basic notations. Let $V$ be a superspace, i.e., a $\mathbb{Z}_{2}$-graded vector space $V=V_{\overline{0}}\oplus V_{\overline{1}}$, where $\{{\overline{0}},\overline{1}\}=\mathbb{Z}_{2}$. For $a\in V$, we say that the element $a$ has parity $p(a)\in\mathbb{Z}_{2}$ if $a\in V_{p(a)}$. A field $a(z)$ is a formal series of the form
$a(z)=\displaystyle\sum_{n\in\mathbb{Z}}a_{(n)}z^{-n-1},\ \ a_{(n)}\in\operatorname{End}(V)$
such that for any $v\in V$, one has $a_{(n)}v=0$ for some $n\gg 0$.
Definition 2.1.
A vertex superalgebra (VSA) refers to a quadruple
$(V,\mathbf{1},T,Y(-,z)),$
where $V=V_{\bar{0}}\oplus V_{\bar{1}}$ is a superspace (state space), $\mathbf{1}\in V_{\overline{0}}$ (vacuum vector), $T\in\operatorname{End}(V)$ (derivation), and
$Y(-,z)\in\operatorname{End}(V)[[z^{\pm}]],\ \ Y(a,z)=:a(z)=\sum_{n\in\mathbb{Z}}a_{(n)}z^{-n-1}$
(the state-field correspondence) satisfying the following axioms:
  • For any $a\in V$, $a(z)$ is a field,
  • (translation coinvariance): $T,a(z)]=\partial a(z)$,
  • (vacuum): $\mathbf{1}(z)=Id_{V}$, $a(z)\mathbf{1}|_{z=0}=a,$
  • (locality): $(z-w)^{N}a(z)b(w)=(-1)^{p(a)p(b)}(z-w)^{N}b(w)a(z)$ for $N\gg 0$.
By abuse of notation, we simply write $V$ for the corresponding VSA. Other representation theoretic concepts (e.g., homomorphism, module,…) are defined in a common way and we omit it (for more detail, see [20]). For a VSA $V$, the parity automorphism $\iota_{V}\in\operatorname{Aut}(V)$ is defined by $\iota_{V}(a)=p(a)a$. Note that $\iota_{V_{1}\otimes V_{2}}=\iota_{V_{1}}\otimes\iota_{V_{2}}$.
Note that from the locality axiom, one can derive the following commutator formula
$a_{(n)}b(z)=b(z)a_{n}+\sum_{i\geq 0}\binom{n}{i}z^{n-i}(a_{(i)}b)(z).$
We call $a_{(0)}$ a zero-mode. By the commutator formula, zero-mode is a derivation:
(2.1) $[a_{(0)},b_{(n)}]=(a_{(0)}b)_{(n)}.$
Definition 2.2.
A VSA $V$ is called a $\tfrac{1}{2}\mathbb{Z}$-graded vertex operator superalgebra (VOSA) if it is $\frac{1}{2}\mathbb{Z}$-graded $V=\oplus_{n\in\frac{1}{2}\mathbb{Z}}V_{n}$ and has a conformal vector $\omega\in V_{2}$ such that $\omega_{(0)}=T$ and the set of operators $\{L_{n}:=\omega_{(n+1)}\}_{n\in\mathbb{Z}}$ satisfy the relation of Virasoro algebra
(2.2) $[L_{m},L_{n}]=(m-n)L_{m+n}+\tfrac{m^{3}-m}{12}\delta_{m+n,0}c_{V},\ \ [L_{n},c_{V}]=0,\ \ (m,n\in\mathbb{Z}).$
[…]
$M=\bigoplus_{\Delta,x_{1},\dots,x_{n}}M_{\Delta,x_{1},\dots,x_{n}},\ \ \dim_{\mathbb{C}}M_{\Delta,x_{1},\dots,x_{n}}<\infty$
for the grading. The character $\operatorname{ch}_{q,z_{1},\dots,z_{n}}M$ (resp. supercharacter $\operatorname{sch}_{q,z_{1},\dots,z_{n}}M$) of $M$ is defined by
[…]
A graded VOSA-module $M$ is CFT-type if there exists $m\in M_{\Delta,x_{1},\dots,x_{n}}$ for some $\Delta,x_{1},\dots,x_{n}$ such that $M_{\Delta,x_{1},\dots,x_{n}}=\mathbb{C}m$ and $M=U(V)m$. Note that a VOSA is always CFT-type because $a=a_{(-1)}\mathbf{1}$. If we do not consider $a^{1},\dots,a^{n}$, we will call the grading conformal grading. After Section 2.3, we only work with conformally-graded VOSAs and their modules. [^8] On the other hand, for example in [14], we consider the grading by $(L_{0},a^{1}_{(0)}:=h_{(0)})$ for $h\in\mathfrak{sl}_{2}\subset\hat{\mathfrak{sl}_{2}}$.
If we allow the exponent of $z$ in field expansion to be a rational number, then one can define the generalized V(O)SA in a similar manner (see [18] for more details). Let us introduce the following two operations that create another VOSA module from a given one.
Definition 2.3.
Let $V$ be a VO(S)A and $M$ be a graded $V$-module.
  1. (1)
    (Contragredient dual) The contragredient dual $M^{\ast}$ of $M$ is defined as the restricted dual of $M$, i.e., $M^{\ast}=\bigoplus_{\Delta,x_{1},\dots,x_{n}}M_{\Delta,x_{1},\dots,x_{n}}^{*}$ with the following VO(S)A-module structure [25]:
    $\langle Y_{M^{\ast}}(v,z)m^{\prime},m\rangle=\langle m^{\prime},Y_{M}(e^{zL_{1}}(-z^{-2})^{L_{0}}v,z^{-1})m\rangle=:\langle m^{\prime},Y_{M}(v,z)^{\dagger}m\rangle,$
    where $v\in V$, $m\in M$, $m^{\prime}\in M^{\ast}$, and $\langle m^{\prime},m\rangle=m^{\prime}(m)$. In other words,
    $v_{(n)}^{\dagger}m=\sum_{\Delta\in\mathbb{Z}}(-1)^{\Delta}\sum_{m\geq 0}(\tfrac{L_{1}^{m}}{m!}v_{\Delta})_{(-n-m-2\Delta-2)}$
    for the conformal decomposition $v=\sum_{\Delta\in\mathbb{Z}}v_{\Delta}$, $v_{\Delta}\in V_{\Delta}$. Direct calculation shows that
    (2.3) $\langle m^{\prime},L_{n}m\rangle=\langle L_{-n}m^{\prime},m\rangle,\ \ (a^{i}_{(0)})^{\dagger}=a^{i}_{(0)}.$
  2. (2)
    (Twisted module and spectral flow twist) For $g\in\operatorname{Aut}(V)$ with finite order $s\in\mathbb{Z}_{\geq 1}$, we can define $g$-twisted modules of $V$, which play an important role in orbifold conformal theory (see [39] for more details). Let us give a procedure called spectral flow twist which deform a given twisted $V$-module. Suppose that there exists $h\in V_{\overline{0}}$, called simple current element, and some $k\in\mathbb{C}$ such that
    $L_{n}h=\delta_{n,0}h\ (n\geq 0),\quad h_{(n)}h=k\delta_{n,1}\mathbf{1},\quad\mathrm{Spec}(h_{(0)})\in\tfrac{1}{s}\mathbb{Z},$
    where $h_{(0)}$ acts semisimplely on $V$ and $\mathrm{Spec}(h_{(0)})$ refers to the set of eigenvalues of $h_{(0)}$. Then one can define Li’s Delta operator [40],
    $\Delta(z):=z^{h_{(0)}}\exp{\sum_{n=1}^{\infty}\tfrac{h_{(n)}}{-n}(-z)^{-n}}.$
    Let $M$ be a $g$-twisted $V$-module. Then $(M,Y_{M}(\Delta(z)\cdot,z))$ is a $gg_{h}$-twisted $V$-module, called spectral flow twist of $M$ and denote $\mathcal{S}_{h}(M)$, where $g_{h}=\exp(2\pi ih_{(0)})$ is an automorphism of $V$ with order $s$ (see [39, Proposition 5.4]). Note that if $M$ is irreducible, then so is $\mathcal{S}_{h}(M)$.
2.2. Simplicity theorem
This subsection will not be used until Section 3 and might be skipped once. In this subsection, we formalize the discussion in [46] to enhance the perspective and flexibility of the discussion.999 However, the authors expect that there should be a more concise way leading to the simplicity theorem in our case (ideally, they would like to include this subsection in Part I). One reason for this belief is that Theorem 1.1(2) has strong similarity to [32, II,5], but in [32, II,5] the simplicity theorem is proved as a direct corollary of the BWB theorem (i.e., no need to go through Kazhdan–Lusztig polynomials, etc.). It is a future work to study and develop the theory of shift system independently of VOA. Throughout this subsection, $(\Lambda,\uparrow,\{V_{\lambda}\}_{\lambda\in\Lambda})$ is a shift system.
Lemma 2.4.
[46, Lemma 2.27, Corollary 3.3] For $\lambda,\lambda^{\prime}\in\Lambda$ satisfying the condition (strong), if
$V_{\lambda}^{\ast}\simeq V_{w_{0}\ast\lambda^{\prime}}(-w_{0}(w_{0}\uparrow\lambda^{\prime})),\ w_{0}\uparrow\lambda-w_{0}(w_{0}\uparrow\lambda^{\prime})=-2\rho$
(where $V_{\lambda}^{\ast}$ is the contragredient dual of $V_{\lambda}$ with the contragredient $B$-action), then we have a natural isomorphism
$H^{0}(G\times_{B}V_{\lambda})\simeq H^{0}(G\times_{B}V_{\lambda^{\prime}})^{\ast}.$
In particular, if $\lambda=\lambda^{\prime}$, then $H^{0}(G\times_{B}V_{\lambda})$ is self-dual.
Proof.
By Theorem 1.1(1) and the Serre duality, we have
$H^{0}(G\times_{B}V_{\lambda})\simeq H^{l(w_{0})}(G\times_{B}V_{w_{0}\ast\lambda}(w_{0}\uparrow\lambda))\simeq H^{l(w_{0})}(G\times_{B}V_{\lambda^{\prime}}^{\ast}(-2\rho)))\simeq H^{0}(G\times_{B}V_{\lambda^{\prime}})^{\ast}$
as $G$-modules. For the naturality discussion, see [46, Section 2]. ∎
Remark 2.5.
In [14, Proposition 6.12], the isomorphisms in Lemma 2.4 hold up to the spectral flow twist $\mathcal{S}^{2}$. However, since $\mathcal{S}^{2}$ commutes with the $B$-action, it does not affect to the discussion and we omit it.
Lemma 2.6.
Let $(\Lambda,\uparrow,\{V_{\lambda}\}_{\lambda\in\Lambda})$ be a shift system. For a fixed $\beta\in P_{+}$ and $\lambda_{0},\lambda_{1}\in\Lambda$ such that the character of $\mathcal{W}_{-\beta+\lambda_{i}}$ is given by (1.20), let us assume the family of graded vector spaces
$\{\mathbb{M}(y,\mu_{\lambda_{i}})=\bigoplus_{\Delta}\mathbb{M}(y,\mu_{\lambda_{i}})_{\Delta}\}_{y\in\hat{W}},\ \ \mathbb{L}(-\beta+\lambda_{i})=\bigoplus_{\Delta}\mathbb{L}(-\beta+\lambda_{i})_{\Delta}\ \ (i=0,1)$
(where $\hat{W}$ is the affine Weyl group of $\hat{\mathfrak{g}}$, $\mu_{\lambda_{i}}$ is an element in $\hat{\mathfrak{h}}^{\ast}$ and denote $y\sim y^{\prime}$ if $y^{\prime}\in Wy$) such that
  1. (1)
    $\operatorname{ch}\mathbb{M}(y,\mu_{\lambda_{i}})=\operatorname{ch}\mathbb{M}(y^{\prime},\mu_{\lambda_{i}})$ iff $y\sim y^{\prime}$, and $\{\operatorname{ch}\mathbb{M}(y,\mu_{\lambda_{i}})\}_{y\in\hat{W}/\sim}$ are linearly independent.
  2. (2)
    $\operatorname{ch}\mathbb{M}(y_{\sigma},\mu_{\lambda_{i}})=\operatorname{ch}V_{\lambda_{i}}^{h=\sigma\circ\beta}$ for some $y_{\sigma}\in\hat{W}$ s.t. $y_{\sigma}\sim y_{\sigma^{\prime}}$ iff $\sigma=\sigma^{\prime}$,
  3. (3)
    $\operatorname{ch}\mathbb{L}(-\beta+\lambda_{i})=\sum_{y\in\hat{W}}a_{\beta,y}\operatorname{ch}\mathbb{M}(y,\mu_{\lambda_{i}})$ for some $a_{\beta,y}\in\mathbb{C}$ (note that $a_{\beta,y}$ is independent of $\lambda_{i}$),
  4. (4)
    $\operatorname{ch}\mathcal{W}_{-\beta+\lambda_{0}}=\operatorname{ch}\mathbb{L}(-\beta+\lambda_{0})$.
Then we have $\operatorname{ch}\mathcal{W}_{-\beta+\lambda_{1}}=\operatorname{ch}\mathbb{L}(-\beta+\lambda_{1})$.
Proof.
By assumption, we have
[…]
By applying (1) and (2) to the equation, we have
$\sum_{y^{\prime}\sim y}a_{\beta,y^{\prime}}=\begin{cases}(-1)^{l(\sigma)}&y\sim y_{\sigma},\\ 0&\text{otherwise}.\end{cases}$
By applying this relation to $\operatorname{ch}\mathcal{W}_{-\beta+\lambda_{1}}$, the assertion is proved. ∎
In the remainder of this subsection, we consider the case Example 1.8 and the following conditions:
  1. (1)
    $W_{Q}:=H^{0}(G\times_{B}V_{0})$ is a generalized CFT-type VOSA and $W_{Q+\lambda}:=H^{0}(G\times_{B}V_{\lambda})$ is a graded $W_{Q}$-module for each $\lambda\in\Lambda$. Furthermore, the natural $G$-action is in $\operatorname{Aut}(W_{Q})$.
  2. (2)
    $W_{P}:=\bigoplus_{\begin{subarray}{c}\mu\in\Lambda\\ \mu_{\bullet}=0\end{subarray}}H^{0}(G\times_{B}V_{\mu})$ is a generalized CFT-type VOSA and $W_{P+\lambda_{\bullet}}:=\bigoplus_{\begin{subarray}{c}\mu\in\Lambda\\ \mu_{\bullet}=\lambda_{\bullet}\end{subarray}}H^{0}(G\times_{B}V_{\mu})$ is a graded $W_{P}$-module for each $\lambda\in\Lambda$. Furthermore, the natural $G$-action is in $\operatorname{Aut}(W_{P})$ and $W_{P}$ is an abelian intertwining algebra with abelian group $P/Q$ (see [43]).
If (2) holds, then so does (1) (but the converse does not in general). If (1) (resp. (2)) holds, then $\mathcal{W}_{0}\simeq W_{Q}^{G}\simeq W_{P}^{G}$ is a vertex operator super subalgebra of $W_{Q}$ and $\mathcal{W}_{-\beta}$ is a $\mathcal{W}_{0}$-module for any $\beta\in P_{+}\cap Q$ (resp. for any $\beta\in P_{+}$) and $\lambda\in\Lambda$. In the same manner as [46, Section 3.2], we obtain the following.
Lemma 2.7.
Let us assume that $\lambda_{\bullet}=0$ satisfies (strong) and (1) (resp. (2)) holds. Then $W_{Q}$ (resp. $W_{P}$) is a simple VOSA (resp. simple generalized VOSA). Furthermore, for each $\beta\in P_{+}\cap Q$ (resp. $\beta\in P_{+}$), $\mathcal{W}_{-\beta}$ is a simple $\mathcal{W}_{0}$-module. If (2) holds, then $W_{Q+\lambda^{\bullet}}$ is a simple $W_{Q}$-module.
Proof.
We will omit the detail of the proof (see [46, Section 3.2]). Let us consider the case (2). By Lemma 2.4, $W_{P}$ is self-dual. Then for the non-degenerate $W_{P}$-invariant bilinear form $\langle\ ,\ \rangle\colon W_{P}\times W_{P}\rightarrow\mathbb{C}$, we have
$1=\langle a^{\ast},a\rangle=\langle a^{\ast}_{(-1)}\mathbf{1},a\rangle=\langle\mathbf{1},(a^{\ast}_{(-1)})^{\dagger}a\rangle,$
where $a\in W_{P}$ and $a^{\ast}\in W_{P}\simeq W_{P}^{\ast}$ by abuse of notation. Since $W_{P}$ is CFT-type, we have $(a^{\ast}_{(-1)})^{\dagger}a=\mathbf{1}$ and thus $W_{P}$ is simple. By [43, Theorem 3.2], $\mathcal{W}_{-\beta}$ is simple as $\mathcal{W}_{0}$-module for any $\beta\in P_{+}$. Finally, by [43, Proposition 2.26], each $W_{Q+\lambda^{\bullet}}$ is simple as $W_{Q}$-module. ∎
In the same manner as [46, Lemma 3.5, 3.22] or [14, Theorem 6.15], we have the following.
Lemma 2.8.
Let us assume (1) and for some $\lambda\in\Lambda$ and any $\alpha\in P_{+}\cap Q$, $\mathcal{W}_{-\alpha+\lambda_{\bullet}}$ is CFT-type as $\mathcal{W}_{0}$-module by a certain $|\lambda_{\bullet}\rangle\in\mathcal{W}_{-\alpha+\lambda_{\bullet}}$. Then $W_{Q+\lambda_{\bullet}}$ is simple as $W_{Q}$-module. Furthermore, if (2) holds and $\mathcal{W}_{-\alpha+\lambda}$ is CFT-type as $\mathcal{W}_{0}$-module by a certain $|\lambda\rangle\in\mathcal{W}_{-\alpha+\lambda}$, then $W_{Q+\lambda}$ is simple as $W_{Q}$-module.
Proof.
We will omit the detail of the proof by the same reason above. Let us consider the later half. Using the Leibniz rule (2.1) repeatedly, $W_{P+\lambda_{\bullet}}$ is generated by $|\lambda_{\bullet}\rangle$ as $W_{P}$-module. In the same manner as Lemma 2.7, $W_{P+\lambda_{\bullet}}$ is simple as $W_{P}$-module (by the same discussion for $W_{Q}$ and $W_{Q+\lambda_{\bullet}}$, the first half of the Lemma is already proved at this point). On the other hand, $W_{Q+\lambda}$ is generetad by some element in $L_{\lambda^{\bullet}}\otimes\mathbb{C}^{\times}|\lambda\rangle$ as $W_{Q}$-module. Since $W_{Q+\lambda}\subseteq W_{P+\lambda_{\bullet}}$ and the $W_{P+\lambda_{\bullet}}$ is simple as $W_{P}$-module, by [19, Corollary 4.2], any element of $W_{Q+\lambda}$ has the form $\sum_{n\in\mathbb{Z}}X^{n}_{(n)}|\lambda\rangle$ for some $\{X^{n}\}_{n\in\mathbb{Z}}\subset W_{P}$. By weight consideration, we can show that $\{X^{n}\}_{n\in\mathbb{Z}}\subset W_{Q}$, and thus $W_{Q+\lambda}$ is simple as $W_{Q}$-module. ∎
Let us explain how these results will be used to the simplicity theorems of (multiplet) W-(super)algebras. We consider the case (1) or (2). Then, by Lemma 2.7, $\mathcal{W}_{-\beta}$ is simple as $\mathcal{W}_{0}$-module. To extend the simplicity of $\mathcal{W}_{-\beta}$ to $\mathcal{W}_{-\beta+\lambda}$ such that $\lambda\in\Lambda$ satisfies (strong), we consider using Lemma 2.6 to $\lambda_{0}=0$ and $\lambda_{1}=\lambda$. Here, $\mathbb{M}(\mu)$ and $\mathbb{L}(\mu)$ represent the Verma module and the irreducible module over $\mathcal{W}_{0}$ with highest weight $\mu$, respectively. As we show later, sometimes $\mathbb{L}(-\beta+\lambda_{i})$ ($i=0,1$) has a ”Kazhdan–Lusztig-type” character formula in Lemma 2.6(3). Since $\mathcal{W}_{-\beta+\lambda_{0}}$ is irreducible, by Lemma 2.6, so is $\mathcal{W}_{-\beta+\lambda_{1}}$. In particular, $\mathcal{W}_{-\beta+\lambda}$ are cyclic. Furthermore, if they are CFT-type, by Lemma 2.8, $W_{Q+\lambda_{\bullet}}$ or $W_{Q+\lambda}$ is also simple as $W_{Q}$-module.
In [45, 46] and [14], the cases where $\mathcal{W}_{0}$ is the principal W-algebra $\mathcal{W}_{k}(\mathfrak{g})$ for simply-laced $\mathfrak{g}$ and the affine VOA $V_{k}(\mathfrak{sl}_{2})$ were considered, respectively, and the corresponding simplicity theorems are proved in the same manner explained above. In Section 3 and Section 4 of the present paper, we consider the cases where $\mathcal{W}_{0}$ is the principal W-(supr)algebra $\mathcal{W}_{k}(\mathfrak{g})$ for non-simply-laced $\mathfrak{g}$ and $\mathfrak{g}=\mathfrak{osp}(1|2r)$, respectively. In the former case, (2) holds and the simplicity theorem will be proved for any $\lambda\in\Lambda$ satisfying (strong), but in the later case, it will be proved only for $\lambda^{\bullet}=0$ case (i.e., assume only (1)) because of some technical difficulty.
2.3. Free field algebras
Let us introduce two VOSAs that will serve as material for shift systems later.
2.3.1. Fermionic VOSA
The Clifford algebra is generated by $\{\phi(n),n\in\frac{1}{2}+\mathbb{Z}\}\cup\{{1}\}$ satisfying the relations
$\{\phi(n),\phi(m)\}=\delta_{n,-m},\quad n,m\in\tfrac{1}{2}+\mathbb{Z}.$
Let $F$ be the module of Clifford algebra generated by $\mathbb{1}$ such that $\phi(n)\mathbb{1}=0$ ($n>0$). Then there is a unique VOSA structure on $F$, with the field and conformal vector
(2.4) $Y(\phi(-\tfrac{1}{2})\mathbb{1},z)=\phi(z)=\sum_{n\in\frac{1}{2}+\mathbb{Z}}\phi(n)z^{-n-\frac{1}{2}},\ \ \omega^{(s)}=\tfrac{1}{2}\phi(-\tfrac{3}{2})\phi(-\tfrac{1}{2})\mathbb{1},$
which gives the central charge $\tfrac{1}{2}$. The conformal weight of $\Delta_{\phi}$ of $\phi(-\tfrac{1}{2})\mathbf{1}$ is $\tfrac{1}{2}$. The (super)characters of $F$ are
$\operatorname{ch}_{q}F=\tfrac{\eta(q)^{2}}{\eta(q^{2})\eta(q^{\frac{1}{2}})},\ \ \operatorname{sch}_{q}F=\tfrac{\eta(q^{\frac{1}{2}})}{\eta(q)},\ \ \operatorname{ch}\iota^{\ast}F=2\tfrac{\eta(q^{2})}{\eta(q)}\ \ \text{($\iota$ is the parity isomorphism)},$
respectively. Moreover, $(\phi(-\frac{1}{2})v^{*},v)=(v^{*},\text{i}\phi(-\frac{1}{2})v)$ for the dual $v^{\ast}\in F^{\ast}$ of $v\in F$.
2.3.2. Lattice VOSAs
Let $L$ be a positive definite integral lattice of rank $r$. We extend the bilinear form on $L$ to $\mathfrak{h}_{L}=\mathbb{C}\otimes_{\mathbb{Z}}L$. We also consider the affinization $\hat{\mathfrak{h}}_{L}=\mathbb{C}[t,t^{-1}]\otimes\mathfrak{h}_{L}\oplus\mathbb{C}c$ of $\mathfrak{h}_{L}$ and its subalgebras $\hat{\mathfrak{h}}_{L}^{\geq 0}=\mathbb{C}[t]\otimes\mathfrak{h}_{L}$, $\hat{\mathfrak{h}}_{L}^{-}=t^{-1}\mathbb{C}[t^{-1}]\otimes\mathfrak{h}_{L}$. For $\mu\in\mathfrak{h}$, the induced representation $M(\mu)$ of $\hat{\mathfrak{h}}_{L}$ is defined by
$M(\mu)=U(\hat{\mathfrak{h}}_{L})\otimes_{U(\hat{\mathfrak{h}}_{L}^{\geq 0}\oplus\mathbb{C}c)}\mathbb{C}_{\mu},$
where $h\in\hat{\mathfrak{h}}_{L}$ acts as $(\mu,h)$ on $\mathbb{C}_{\mu}$. Then there exists a simple V(S)A structure (see [33]) on
$V_{L}=M(0)\otimes\mathbb{C}[L]\simeq\bigoplus_{\mu\in L}M(\mu)$
with the vacuum vector $\mathbf{1}=1\otimes 1$, such that $Y(ht^{-1}\otimes 1,z)=h(z)=\sum_{n\in\mathbb{Z}}h_{(n)}z^{-n-1}$, $h_{(n)}=h\otimes t^{n}$. The group algebra of $\mathbb{C}[L]$ is generated by $e^{\mu}$, $\mu\in{L}$. The vertex operator of $Y(e^{\mu},z)$ is defined as
$Y(e^{\mu},z)=e^{\mu}z^{\mu_{(0)}}\exp\left(-\sum_{j<0}\tfrac{z^{-j}}{j}\mu_{(j)}\right)\exp\left(-\sum_{j>0}\tfrac{z^{-j}}{j}\mu_{(j)}\right)\epsilon(\mu,\cdot),$
where $e^{\mu}(e^{\lambda})=e^{\mu+\lambda}$ for $\lambda\in\mathbb{C}[{L}]$, and $\epsilon(\cdot,\cdot)$ is the $2$-cocycle satisfying [33, (5.4.14)]. The parity of the element $v\otimes e^{\mu}$ is $|\mu|^{2}\mod{2}$. One can define a (shifted) conformal vector
(2.5) $\omega_{\mathrm{sh}}=\omega_{\mathrm{st}}+\gamma_{(-2)}\mathbf{1},\ \ \omega_{\mathrm{st}}=\tfrac{1}{2}\sum_{i=1}^{r}v_{i(-1)}v_{i}^{\ast},\ \ \gamma\in\mathfrak{h}_{L}$
with central charge is $r-12(\gamma,\gamma)$, where $\{v^{i}\}_{i=1}^{\ell}$ and $\{v_{i}^{\ast}\}_{i=1}^{r}$ are bases of $L$ and $L^{\ast}$, respectively, such that $(v_{i},v_{j}^{\ast})=\delta_{i,j}$. The dual lattice $L^{\ast}$ of $L$ is defined by $L^{\ast}:=\{\beta\in\mathbb{Q}\otimes L\;|\;(\alpha,\beta)\in\mathbb{Z}\,\,\text{for all}\,\,\alpha\in L\}$. Then the lattice VO(S)A has finitely many irreducible modules parametrized by $L^{\ast}/L$ [41]; these modules are
$V_{L+\lambda}:=M(0)\otimes e^{\lambda}\mathbb{C}[L],\ \ (\lambda\in L^{\ast}/L)$
up to isomorphism. For $\mu\in\lambda+L$, the conformal weight $\Delta_{\mu}$ of $e^{\mu}$ under the action of $(\omega_{\mathrm{sh}})_{(1)}=L_{0}$ is
(2.6) $\Delta_{\mu}=\tfrac{1}{2}|\mu-\gamma|^{2}+\tfrac{c-r}{24}=\tfrac{1}{2}|\mu|^{2}-(\mu,\gamma).$
According to [25, Proposition 5.3.2], the contragredient dual $V_{L+\lambda}^{\ast}$ of $V_{L+\lambda}$ is $V_{L+\lambda^{\prime}}$, for some $\lambda^{\prime}\in L^{\ast}/L$.
2.3.3. Ramond sector
Let us consider the setup in the last two sections and Definition 2.3(2). For
$R=\tfrac{1}{2}\sum_{i\in I_{\mathrm{odd}}}v_{i}^{\ast},\ \ I_{\mathrm{odd}}=\{1\leq i\leq r\ |\ \text{$|v_{i}|^{2}$ is odd}\},$
we can easily check that $R$ is a simple current element and $g_{R}=\operatorname{exp}(2\pi iR_{(0)})$ gives the parity automorphism $\iota_{V_{L}}$ of $V_{L}$. Let $M$ be a $1\otimes\iota_{F}$-twisted $V_{L}\otimes F$-module. Note that $\iota_{V_{L}}\otimes 1\in\operatorname{Aut}(V_{L}\otimes F)$. The Ramond sector of $M$ is defined by the $\iota_{V_{L}\otimes F}$-twisted $V_{L}\otimes F$-module $\mathcal{S}_{R}(M)$.
2.4. Construction of shift systems
Let us consider step-by-step if we can build a shift system (Definition 1.6) using the two VOSA(-modules), free fermion $F$ and lattice VOSA-module $V_{L+\lambda}$ with conformal vectors (2.4) and (2.5), respectively, introduced above. To consider $V_{L+\lambda}$ as a weight $B$-module, it is natural to consider the case $L=xQ$ for some $x\in\mathbb{C}^{\times}$ and the $\mathfrak{h}$-action
(2.7) $h_{i}=\lceil-\tfrac{1}{x}\alpha_{i(0)}^{\vee}\rceil.$
In other words, we consider the Fock space $M(-x\mu)$ as the Cartan weight space with weight $\lceil\mu\rceil$, where $\lceil\mu\rceil\in P$ is uniquely determined by $\lceil(\mu,\alpha_{i}^{\vee})\rceil=(\lceil\mu\rceil,\alpha_{i}^{\vee})$ for $i\in\Pi$. We consider the case where $L=xQ$ is positive integral (see Section 2.3.2), which is equivalent to the cases $x=\sqrt{p}$ for some $p\in\mathbb{Z}_{\geq 1}$ such that
  1. (1)
    $p=r^{\vee}m$ (where $r^{\vee}$ is the lacing number of $\mathfrak{g}$ and $m\in\mathbb{Z}_{\geq 1}$), or
  2. (2)
    p=2m1p=2m-1, $m\in\mathbb{Z}_{\geq 1}$ and $\mathfrak{g}=B_{r}$.
Note that in the case (1), $L=\sqrt{p}Q$ is even, but in the case (2), $\sqrt{p}Q$ is odd (because $|\sqrt{p}\alpha_{r}|^{2}$ is odd). Also, since $\sqrt{p}Q$ is an integral, $\sqrt{p}Q\subseteq(\sqrt{p}Q)^{\ast}$ and thus we can consider $(\sqrt{p}Q)^{\ast}/\sqrt{p}Q$. For the moment, we will not fix the representative $\Lambda$ of $(\sqrt{p}Q)^{\ast}/\sqrt{p}Q\simeq\tfrac{1}{p}Q^{\ast}/Q$, and $\lambda$ will simply represent an element of $\tfrac{1}{p}Q^{\ast}$ and we use the letter $V_{\sqrt{p}(Q+\lambda)}$ for the corresponding irreducible $V_{\sqrt{p}Q}$-module. The representative $\Lambda$ will be determined later from considerations on the choice of conformal weight vector.
Since the $\mathfrak{h}$-action is given by (2.7), it is natural to consider whether the zero-modes
$e^{\sqrt{p}\alpha_{i}}_{(0)}\colon V_{\sqrt{p}(Q+\lambda)}^{h=\lceil\mu\rceil}=M(-\sqrt{p}\mu)\rightarrow M(-\sqrt{p}(\mu-\alpha_{i}))=V_{\sqrt{p}(Q+\lambda)}^{h=\lceil\mu-\alpha_{i}\rceil=\lceil\mu\rceil-\alpha_{i}}$
define a $B$-action of a shift system. For a $B$-action to be compatible with the structure of VOSA, it should be even, but in the case (2), $e^{\sqrt{p}\alpha_{r}}_{(0)}$ is odd as we checked above. To avoid this problem, in the case of (2), let us consider $V_{\sqrt{p}(Q+\lambda)}\otimes F$ and $(e^{\sqrt{p}\alpha_{r}}\otimes\phi)_{(0)}$ instead of $V_{\sqrt{p}Q}$ and $e^{\sqrt{p}\alpha_{r}}_{(0)}$, respectively. Then since $\phi\in F$ is odd, $(e^{\sqrt{p}\alpha_{r}}\otimes\phi)_{(0)}$ is even. From now on we consider the following two cases: For case (1) and (2),
(2.8) $V_{\lambda}=V_{\sqrt{p}(Q+\lambda)},\ \ \omega=\omega_{\mathrm{sh}},\ \ h_{i}=\lceil-\tfrac{1}{\sqrt{p}}\alpha_{i(0)}^{\vee}\rceil,\ \ f_{i}=e^{\sqrt{p}\alpha_{i}}_{(0)}\ \ (1\leq i\leq r),$
(2.9) $V_{\lambda}=V_{\sqrt{p}(Q+\lambda)}\otimes F,\ \ \omega=\omega_{\mathrm{sh}}+\omega^{(s)},\ \ h_{i}=\lceil-\tfrac{1}{\sqrt{p}}\alpha_{i(0)}^{\vee}\rceil,\ \ f_{i}=\begin{cases}e^{\sqrt{p}\alpha_{i}}_{(0)}&(1\leq i\leq r-1),\\ (e^{\sqrt{p}\alpha_{r}}\otimes\phi)_{(0)}&(i=r),\end{cases}$
respectively. Recall that in Definition 1.6(3a), we need a grading of $V_{\lambda}$ which decomposes $V_{\lambda}$ into finite dimensional weight $B$-modules. Here we shall employ the conformal grading by $\omega$ above. Then, in order for the operators $\{f_{i},h_{i}\}_{i\in\Pi}$ to define the $B$-action in the shift system, in addition to the Serre relation and integrability, the conformal weights must be preserved [^10] Note that compatibility with the entire Virasoro action is not necessarily required.. In each cases (1), (2), let us determine such a conformal vector $\omega$ and describe the shift system naturally derived from the choice of $\omega$. In particular, we will give the pair $(\Lambda,\uparrow)$ as special cases of Example 1.8, and thus Definition 1.6(1),(2) are satisfied.
2.4.1. The case (1)
By (2.6), the operators $f_{i}$ ($i\in I$) preserve the conformal weight iff
(2.10) $\omega=$ $\tfrac{1}{2}\sum_{i\in\Pi}\alpha_{i(-1)}\alpha_{i}^{\ast}+\sqrt{p}(\rho-\tfrac{1}{p}\rho^{\vee})_{(-2)}\mathbf{1}$
and thus we consider this conformal vector. On the other hand, under the choice (2.10), another solution of $\Delta_{e^{x\alpha_{i}}_{(0)}}=0$ is $x=-\tfrac{2}{\sqrt{p}(\alpha_{i},\alpha_{i})}$, namely the case $x\alpha_{i}=-\tfrac{1}{\sqrt{p}}\alpha_{i}^{\vee}$. By [49], if $(p\lambda+\rho^{\vee},\alpha_{i})\equiv s\ (\operatorname{mod}\ p)$ for $1\leq s\leq p-1$, then we have the non-zero (conformal weight-preserving) screening operators
(2.11) $Q_{i,\lambda}=\int_{[\Gamma]}e^{-\tfrac{1}{\sqrt{p}}\alpha_{i}^{\vee}}(z_{1})\cdots e^{-\tfrac{1}{\sqrt{p}}\alpha_{i}^{\vee}}(z_{s})dz_{1}...d_{z_{s}}\colon V_{\lambda}\rightarrow V_{\lambda-\tfrac{s}{p}\alpha_{i}^{\vee}},$
where $\Gamma]$ is the cycle defined in [44, Section 2.3]. To use (2.11) as short screening operators of a shift system, we have to regard $\lambda-\tfrac{s}{p}\alpha_{i}^{\vee}$ as $\sigma_{i}\ast\lambda$ for some $W$-action $\ast$. If we choose $x=\tfrac{1}{p}\rho^{\vee}$ in Example 1.8, then the representative $\Lambda$ of $\tfrac{1}{p}Q^{\ast}/Q$ is given by
(2.12) $\Lambda=\{-\lambda^{\bullet}+\lambda_{\bullet}\ |\ \lambda^{\bullet}\in P_{\mathrm{min}},\ \lambda_{\bullet}=\sum_{i\in\Pi}\tfrac{\lambda_{i}-1}{p}\alpha_{i}^{\ast},\ 1\leq\lambda_{i}\leq r_{i}^{\vee}m\},\ r_{i}^{\vee}=\begin{cases}r^{\vee}&(\text{$\alpha_{i}$ is long}),\\ 1&(\text{$\alpha_{i}$ is short}),\end{cases}$
where $P_{\mathrm{min}}$ is the set of minuscule weights, and has the $W$-action $\ast$ and the shift map $\uparrow$ induced from
(2.13) $\sigma\ast\mu=\sigma(\mu+\tfrac{1}{p}\rho^{\vee})-\tfrac{1}{p}\rho^{\vee}.$
Then under the $W$-action, the short screening operator (2.11) is described as $Q_{i,\lambda}\colon V_{\lambda}\rightarrow V_{\sigma_{i}\ast\lambda}(\sigma_{i}\uparrow\lambda)$.
2.4.2. The case (2)
By (2.6) and $\Delta_{\phi}=\tfrac{1}{2}$, the operators $f_{i}$ ($i\in I$) preserve the conformal weight iff
(2.14) $\omega=\omega^{(s)}+\tfrac{1}{2}\sum_{i\in I}\alpha_{i(-1)}\alpha_{i}^{\ast}+\sqrt{p}(1-\tfrac{1}{p})\rho_{(-2)}\mathbf{1}$
and thus we consider this conformal vector (note that the $\omega_{\mathrm{sh}}$-part is different from (2.10)). On the other hand, under the choice (2.14), another solution of $\Delta_{e^{x\alpha_{i}}_{(0)}}=0$ ($i\not=r$) and $\Delta_{(e^{x\alpha_{r}}\otimes\phi)_{(0)}}=0$ is $x=-\tfrac{1}{\sqrt{p}}$. In the same manner as above, if $(p\lambda+\rho,\alpha_{i}^{\vee})\equiv s\ (\operatorname{mod}\ p)$ for $1\leq s\leq p-1$, then we have the non-zero (conformal weight-preserving) screening operators (see also [31])
(2.15) $Q_{i,\lambda}=\begin{cases}\int_{[\Gamma]}e^{-\frac{1}{\sqrt{p}}\alpha_{i}}(z_{1})\cdots e^{-\frac{1}{\sqrt{p}}\alpha_{i}}(z_{s})dz_{1}...d_{z_{s}}&(i\not=r),\\ \int_{[\Gamma]}(e^{-\frac{1}{\sqrt{p}}\alpha_{r}}\otimes\phi)(z_{1})\cdots(e^{-\frac{1}{\sqrt{p}}\alpha_{r}}\otimes\phi)(z_{s})dz&(i=r),\end{cases}\colon V_{\lambda}\rightarrow V_{\lambda-\frac{s}{p}\alpha_{i}}.$
If we choose $x=\tfrac{1}{p}\rho$ in Example 1.8, then the representative $\Lambda$ of $\tfrac{1}{p}Q^{\ast}/Q$ is given by
(2.16) $\Lambda=\{-\lambda^{\bullet}+\lambda_{\bullet}\ |\ \lambda^{\bullet}\in P_{\mathrm{min}},\ \lambda_{\bullet}=\sum_{i\in I}\tfrac{\lambda_{i}-1}{p}\varpi_{i},\ 1\leq\lambda_{i}\leq p,\ \ \lambda_{r}+(\lambda^{\bullet},\alpha_{r}^{\vee})\equiv 1\ (\operatorname{mod}\ 2)\},$
and has the $W$-action $\ast$ and the shift map $\uparrow$ induced from
(2.17) $\sigma\ast\mu=\sigma(\mu+\tfrac{1}{p}\rho)-\tfrac{1}{p}\rho.$
Then under the $W$-action, the short screening operator (2.15) is described as $Q_{i,\lambda}\colon V_{\lambda}\rightarrow V_{\sigma_{i}\ast\lambda}(\sigma_{i}\uparrow\lambda)$.
The following lemma is proved in the same manner as [45, Appendix], so we will omit the proof.
Lemma 2.9.
[45, Appendix]
  1. (1)
    For the $W$-module $\Lambda$ defined by (2.12), (2.13), we have
[…]
    In particular, $m\geq{}^{L}h_{\mathfrak{g}}^{\vee}-1$. Furthermore, for such a $\lambda\in\Lambda$, we have $w_{0}\uparrow\lambda=-\rho^{\vee}$.
  2. (2)
    For the $W$-module $\Lambda$ defined by (2.16), (2.17), we have
[…]
    In particular, $m\geq r$. Furthermore, for such a $\lambda\in\Lambda$, we have $w_{0}\uparrow\lambda=-\rho$.
2.4.3. Verification of Definition 1.6(3)
Let us prove the remaining conditions of the shift system. Since the proofs are similar, we will treat both cases (1) and (2) simultaneously.
Lemma 2.10.
The operators $\{f_{i},h_{i}\}_{i\in I}$ give rise to an integrable $\mathfrak{b}$-action on $V_{\lambda}$.
Proof.
The proof is based on the evaluation of conformal weights. Let us check the Serre relation
$(\operatorname{ad}f_{i})^{1-c_{ij}}f_{j}=0$
($h_{i,\lambda},f_{j}]=-c_{ij}f_{j}$ is clear). Set $f_{i}=x_{(0)}$ and $f_{j}=y_{(0)}$. By (2.1), we have $(\operatorname{ad}f_{i})^{1-c_{ij}}f_{j}=(f_{i}^{1-c_{ij}}y)_{(0)}$ and $f_{i}^{1-c_{ij}}y\in V_{\lambda}^{h=-((1-c_{ij})\alpha_{i}+\alpha_{j})}$. Since $\Delta_{f_{i}}=\Delta_{f_{j}}=0$, we have $\Delta_{f_{i}^{1-c_{ij}}y}=\Delta_{y}=1$. On the other hand, by (2.6) and (2.10), the minimal conformal weight of $V_{\lambda}^{h=-((1-c_{ij})\alpha_{i}+\alpha_{j})}$ is given by
[…]
Since this is greater than $1$, we have $(\operatorname{ad}f_{i})^{1-c_{ij}}f_{j}=0$. We show that the $\mathfrak{b}$-action is integrable. The case of $h_{i}$ is clear and we consider the case of $f_{i}$. For $\beta\in P$, pick any homogeneous element $A$ in $V_{\lambda}^{h=\beta}$. Then we have $f_{i}^{N}A\in V_{\lambda}^{h=\beta-N\alpha_{i}}$ for $N\in\mathbb{N}$. The conformal weight of every nonzero element in $V_{\lambda}^{h=\beta-N\alpha_{i}}$ is greater than or equal to $\Delta_{\sqrt{p}(\beta+N\alpha_{i})+\lambda}$. By (2.6), $\Delta_{\sqrt{p}(\beta+N\alpha_{i})+\lambda}$ is a quadratic function in $N$ with leading coefficient $p(\alpha_{i},\alpha_{i})>0$. Meanwhile, the conformal weight $f_{i}^{N}A$ will stay the same since $f_{i}$ preserves the conformal weight. Thus, if $N$ is big enough such that $\Delta_{\sqrt{p}(\beta+N\alpha_{i})+\lambda}>\Delta_{A}$, then $f_{i}^{N}A=0$. ∎
From now on, we consider the $B$-action on $V_{\lambda}$ in Lemma 2.10.
Lemma 2.11.
The short screening operators $Q_{j,\lambda}\colon V_{\lambda}\rightarrow V_{\sigma_{j}\ast\lambda}(\sigma_{j}\uparrow\lambda)$ are $B$-module homomorphisms.
Proof.
It suffices to check that $f_{i},Q_{j,\lambda}]=0$. First, we consider the case where $s=1$. By calculation, we have $e^{\sqrt{p}\alpha_{i}}_{(r)}e^{-\frac{1}{\sqrt{p}}\alpha_{j}}=0$ for $r\geq 0$, $i\neq j$. Thus, we have $f_{i},{Q}_{j,\lambda}]=0$ for $i\neq j$. When $i\neq n$, we have
(2.18) $e^{\sqrt{p}\alpha_{i}}_{(0)}e^{-\frac{1}{\sqrt{p}}\alpha_{i}}=\sqrt{p}(\alpha_{i})_{(-1)}e^{(\sqrt{p}-\frac{1}{\sqrt{p}})\alpha_{i}}\mathbb{1}=\tfrac{p}{p-1}(L_{-1}e^{(\sqrt{p}-\frac{1}{\sqrt{p}})\alpha_{i}})\mathbb{1}.$
Similarly, when $i=n$, we have
(2.19) $e^{\sqrt{p}\alpha_{n}}_{(-1)}e^{-\frac{1}{\sqrt{p}}\alpha_{n}}=\tfrac{p}{p-1}(L_{-1}e^{(\sqrt{p}-\frac{1}{\sqrt{p}})\alpha_{n}})\mathbb{1}.$
Using commutator formula, (2.18) and (2.19), we have
(2.20) $[f_{i},e^{-\frac{1}{\sqrt{p}}\alpha_{j}}(z)]=0,\quad\text{for}\;\;i\neq j,$
(2.21) $[f_{i},e^{-\frac{1}{\sqrt{p}}\alpha_{i}}(z)]=\tfrac{p}{p-1}\tfrac{d}{dz}e^{-\frac{1}{\sqrt{p}}\alpha_{i}}(z),\quad\text{for}\;\;i\neq n,$
(2.22) $[f_{n},(e^{-\frac{1}{\sqrt{p}}\alpha_{n}}\otimes\phi(-\tfrac{1}{2}))(z)]=\tfrac{p}{p-1}\tfrac{d}{dz}e^{-\frac{1}{\sqrt{p}}\alpha_{n}}(z).$
Note $(L_{-1}u)_{(0)}=0$ in every generalized vertex operator algebra. we conclude that $f_{i},{Q}_{j,\lambda}]=0$ for $1\leq i\leq n$. Now let us consider the case $s\geq 2$. Using (2.20), we have
$[f_{i},{Q}_{j,\lambda}]=0\quad\text{for}\;\;i\neq j,$
$[f_{i},{Q}_{i,\lambda}]=\tfrac{p}{p-1}\int_{[\Gamma_{s_{i}+1}]}\tfrac{d}{dz}Q^{i}(z_{1},...,z_{s_{j}})dz_{1}...dz_{s_{j}},$
where $\tfrac{d}{dz}$ is the total derivation of $Q^{i}(z_{i},...,z_{s_{i}})$. The result follows. ∎
Lemma 2.12.
For $i\in\Pi$ and $\lambda\in\Lambda^{\sigma_{i}}$, we have the short exact sequence of $B$-modules
$0\rightarrow\ker Q_{i,\lambda}\rightarrow V_{\lambda}\rightarrow\ker Q_{i,\sigma\ast\lambda}(\sigma_{i}\uparrow\lambda)\rightarrow 0.$
Furthermore, $\ker Q_{i,\lambda}$ and $\ker Q_{i,\sigma\ast\lambda}$ are the (maximal) $P_{i}$-submodules of $V_{\lambda}$ and $V_{\sigma_{i}\ast\lambda}$, respectively. If $\lambda\in\Lambda^{\sigma_{i}}$, then $V_{\lambda}$ has the $P_{i}$-module structure.
Proof.
Let $M(0)^{i,\perp}$ and $M(0)^{i}$ be the rank r1r-1 and rank $1$ Heisenberg VOAs generated by $\{\varpi_{j}|j\not=i\}$ and $\alpha_{i}$, respectively. Then we have the orthogonal decomposition
$V_{\lambda}\simeq$ $M(0)\otimes e^{\sqrt{p}\lambda}\mathbb{C}[\sqrt{p}Q]\ (\otimes F)$
$\simeq$ $(M(0)^{i,\perp}\oplus M(0)^{i})\otimes e^{\sqrt{p}\lambda}\mathbb{C}[\sqrt{p}Q]\ (\otimes F)$
$\simeq$ $M(0)^{i,\perp}\otimes(\bigoplus_{{\alpha}\in Q/\sim}M(0)^{i}\otimes e^{\sqrt{p}(\alpha+\lambda)}\mathbb{C}[\sqrt{p}\mathbb{Z}\alpha_{i}]\ (\otimes F))$
$=:$ $M(0)^{i,\perp}\otimes(\bigoplus_{{\alpha}\in Q/\sim}V_{\alpha+\lambda}^{i})$
of $V_{\lambda}$, where $\alpha\sim\beta$ iff $\beta\in\alpha+\mathbb{Z}\alpha_{i}$ and the term $(\otimes F)$ is omitted in the case 1. Since $f_{i}$ and $Q_{i,\lambda}$ commutes with the $M(0)^{i,\perp}$-action, we have the orthogonal decompositions
$\ker Q_{i,\lambda}\simeq M(0)^{i,\perp}\otimes(\bigoplus_{{\alpha}\in Q/\sim}\ker Q_{i,\lambda}|_{V_{\alpha+\lambda}^{i}}),\ \ \ker Q_{i,\sigma_{i}\ast\lambda}\simeq M(0)^{i,\perp}\otimes(\bigoplus_{{\alpha}\in Q/\sim}\ker Q_{i,\sigma_{i}\ast\lambda}|_{V_{\alpha+\sigma_{i}\ast\lambda}^{i}})$
of $\ker Q_{i,\lambda}$ and $\ker Q_{i,\sigma_{i}\ast\lambda}$, respectively. If we can show that $\ker Q_{i,\lambda}|_{V_{\alpha+\lambda}^{i}}$ and $\ker Q_{i,\sigma_{i}\ast\lambda}|_{V_{\alpha+\sigma_{i}\ast\lambda}^{i}}$ have the $SL_{2}^{i}$-module structures and the short exact sequence
$0\rightarrow\ker Q_{i,\lambda}|_{V_{\alpha+\lambda}^{i}}\rightarrow V_{\alpha+\lambda}^{i}\rightarrow\ker Q_{i,\sigma_{i}\ast\lambda}|_{V_{\alpha+\sigma_{i}\ast\lambda}^{i}}(\sigma_{i}\uparrow\lambda)\rightarrow 0$
of $B^{i}$-modules, then by Lemma 1.4 and the orthogonal decompositions above, the first-half assertion follows. Therefore, it suffices to consider the cases of rank $1$. These cases have already been studied, and from the choice of $\Lambda$ and the definition of $l$, the following can be directly shown.
  • In the cases where $\alpha_{i}$ is long, the assertion follows from [7, Theorem 1.1,1.2] for $p=r^{\vee}m$.
[…]
Similarly, the last assertion follows from [7, Theorem 1.3] and [9, Theorem 6.3], respectively. ∎
2.4.4. Ramond sector of case (2)
Let us recall the notation in Section 2.3.3. In the case (2), the simple current element $R$ is given by $R=\tfrac{\pi i}{\sqrt{2m-1}}\varpi_{r}$. Under the spectral flow twist $\mathcal{S}_{R}$, the conformal grading is deformed as
(2.23) $\begin{split}L_{0}&=\operatorname{Res}_{z}z^{2}Y(\Delta(\tfrac{\varpi_{r}}{\sqrt{2m-1}},z)\omega_{\mathrm{sh}}+e^{\Delta_{x}}\omega^{(s)},z)\\ &=\operatorname{Res}_{z}z^{2}Y(\omega_{\mathrm{sh}}+\omega^{(s)},z)+\tfrac{1}{16}+\operatorname{Res}_{z}z^{2}Y((\tfrac{(\varpi_{r})_{(1)}}{\sqrt{2m-1}}+\tfrac{1}{2}(\tfrac{(\varpi_{r})_{(1)}(\varpi_{r})_{(1)}-\sqrt{2m-1}(\varpi_{r})_{(2)}}{2m-1}))\omega_{\mathrm{sh}},z)\\ &=:\operatorname{Res}_{z}z^{2}Y(\omega_{\mathrm{sh}}+\omega^{(s)},z)+\tfrac{1}{16}+A(\alpha_{r})_{(0)}+B(\alpha_{r-1})_{(0)}+C,\end{split}$
where $\Delta_{x}$ is defined in [8], and $A$, $B$, $C$ are some constants. In the Ramond sector $\mathcal{S}_{R}V_{\lambda}$, we have
(2.24) $\Delta_{\mu}=\tfrac{1}{2}|\mu|^{2}-\sqrt{2m-1}(1-\tfrac{1}{2m-1})(\mu,\rho)+A(\alpha_{r},\mu)+B(\alpha_{r-1},\mu)+C+\tfrac{1}{16}$
Since a spectral flow twist is invertible, under the appropriate changes, the Ramond sector $\{\mathcal{S}_{R}V_{\lambda}\}_{\lambda\in\Lambda}$ of case (2) also defines a shift system. In conclusion, we obtain the following theorem.
Theorem 2.13.
Let $\mathfrak{g}$ be a finite-dimensional simple Lie algebra and $p\in\mathbb{Z}_{\geq 1}$.
  1. (1)
    If $p=r^{\vee}m$ for $m\in\mathbb{Z}_{\geq 1}$, then we have a shift system $(\Lambda,\uparrow,\{V_{\lambda}\}_{\lambda\in\Lambda})$, where
    • The $W$-module $\Lambda$ is (2.12), where the $W$-action is induced from (2.13).
    • The shift map $\uparrow$ is defined by $\sigma\uparrow\lambda=\sigma\ast\lambda_{\bullet}-(\sigma\ast\lambda)_{\bullet}\in P$.
    • The weight $B$-module $V_{\lambda}$ is (2.8). The grading is defined by the conformal grading of (2.10). The short screening operator $Q_{i,\lambda}$ is given by (2.11).
  2. (2)
    If $\mathfrak{g}=B_{r}$ and p=2m1p=2m-1 for $m\in\mathbb{Z}_{\geq 1}$, then we have a shift system $(\Lambda,\uparrow,\{V_{\lambda}\}_{\lambda\in\Lambda})$, where
    • The $W$-module $\Lambda$ is (2.16), where the $W$-action is induced from (2.17).
    • The shift map $\uparrow$ is defined by $\sigma\uparrow\lambda=\sigma\ast\lambda_{\bullet}-(\sigma\ast\lambda)_{\bullet}\in P$.
    • The weight $B$-module $V_{\lambda}$ is (2.9). The grading is defined by the conformal grading of (2.14). The short screening operator $Q_{i,\lambda}$ is given by (2.15).
    Furthermore, even if $V_{\lambda}$ is replaced by the Ramond sector $\mathcal{S}_{R}V_{\lambda}$, $(\Lambda,\uparrow,\{\mathcal{S}_{R}V_{\lambda}\}_{\lambda\in\Lambda})$ is a shift system.
Finally, the following is proved in the same manner as [46, Lemma 2.18, 2.27] so we omit the proof.
Lemma 2.14.
In Theorem 2.13(1) (resp. (2)), $V_{\lambda}^{\ast}(-2\rho)$ is isomorphic to $V_{w_{0}\ast\lambda^{\prime}}(-\rho^{\vee})$ (resp. $V_{w_{0}\ast\lambda^{\prime}}(-\rho)$) as $B$- and $V_{0}$-modules, where $\lambda^{\prime}$ is the representative of $w_{0}(\lambda)$. In particular, for $0\leq n\leq l(w_{0})$, we have
$H^{n}(G\times_{B}V_{\lambda}^{\ast}(-2\rho))\simeq H^{n}(G\times_{B}V_{w_{0}\ast\lambda^{\prime}}(-\rho^{\vee}))\ \ \text{(resp. $H^{n}(G\times_{B}V_{\lambda}^{\ast}(-2\rho))\simeq H^{n}(G\times_{B}V_{w_{0}\ast\lambda^{\prime}}(-\rho)))$}$
as $G$- and $H^{0}(G\times_{B}V_{0})$-modules.
3. Main results I
In this section, by applying Theorem 1.1 to the shift systems in Theorem 2.13, we show that the main results of [45] and part of [46] also hold for our cases. After stating the definitions and properties common to the two cases of Theorem 2.13, we will examine each case in Section 3.1 and 3.2, respectively.
Definition 3.1.
The multiplet WW-(super)algebra is defined by the vertex operator (super) subalgebra
$W_{\sqrt{p}Q}=\bigcap_{i\in I}\ker Q_{i,0}|_{V_{0}}$
of $V_{0}$. For each $\lambda\in\Lambda$, $\bigcap_{i\in I}\ker Q_{i,\lambda}|_{V_{\lambda}}$ is a $W_{\sqrt{p}Q}$-submodule of $W_{\lambda}$.
On the other hand, in the same manner as [46, Corollary 2.21], the 0-th sheaf cohomology
$H^{0}(G\times_{B}V_{0})$
of the sheaf associated with the $G$-equivariant vector bundle $G\times_{B}V_{0}$ over the flag variety $G/B$ has the VO(S)A structure induced from $V_{0}$, and the $n$-th sheaf cohomology $H^{n}(G\times_{B}V_{\lambda}(\mu))$ is an $H^{0}(G\times_{B}V_{0})$-module (for more detail, see [46, Section 2.1] and the discussion just before [46, Corollary 2.21]). Denote
(3.1) $H^{0}(G\times_{B}V_{\lambda})\simeq\bigoplus_{\alpha\in P_{+}\cap Q}L_{\alpha+\lambda^{\bullet}}\otimes\mathcal{W}_{-\alpha+\lambda}$
the decomposition of $H^{0}(G\times_{B}V_{\lambda})$ as $G$-module, where $L_{\beta}$ is the irreducible $\mathfrak{g}$-module with highest weight $\beta\in P_{+}$ and $\mathcal{W}_{-\alpha+\lambda}$ is the multiplicity of a weight vector of $L_{\alpha+\lambda^{\bullet}}$. In our case,
$H^{0}(G\times_{B}V_{0})^{G}=V_{0}^{B}=\bigcap_{i\in I}\ker f_{i}|_{V_{0}^{h=0}}=\mathcal{W}_{0}$
is a vertex operator (super) subalgebra of $H^{0}(G\times_{B}V_{0})$ and each $\mathcal{W}_{-\alpha+\lambda}$ is a $\mathcal{W}_{0}$-module.
Let us use the same notation and setup in Theorem 2.13. The proofs are basically the same as in [45, 46], so we will only outline them. By Theorem 1.1 and Theorem 2.13 we obtain the following.
Theorem 3.2.
Let $(\Lambda,\uparrow,\{V_{\lambda}\}_{\lambda\in\Lambda})$ be the shift system in Theorem 2.13(1)(resp. (2)).
  1. (1)
    The evaluation map
    $\operatorname{ev}\colon H^{0}(G\times_{B}V_{\lambda})\rightarrow\bigcap_{i\in I}\ker Q_{i,\lambda},\ \ s\mapsto s(\operatorname{id}_{G/B})$
    is injective, and is isomorphic iff $\lambda\in\Lambda$ satisfies the condition (weak). In particular, for any $p\in\mathbb{Z}_{\geq 2}$, we have the isomorphism of VOAs $H^{0}(G\times_{B}V_{\sqrt{p}Q})\simeq W_{\sqrt{p}Q}$.
  2. (2)
    For $\lambda\in\Lambda$ such that $(p\lambda_{\bullet}+\rho^{\vee},{}^{L}\theta)\leq p$ (resp. $(p\lambda_{\bullet}+\rho,{}^{L}\theta)\leq p$), we have a natural $G$- and $W_{\sqrt{p}Q}$-module isomorphism
    $H^{n}(G\times_{B}V_{\lambda})\simeq H^{n+l(w_{0})}(G\times_{B}V_{w_{0}\ast\lambda}(-\rho^{\vee}))$
    $(\text{resp.}\quad H^{n}(G\times_{B}V_{\lambda})\simeq H^{n+l(w_{0})}(G\times_{B}V_{w_{0}\ast\lambda}(-\rho))).$
    In particular, $H^{n>0}(G\times_{B}V_{\lambda})\simeq 0$ and $H^{0}(G\times_{B}V_{\lambda})\simeq H^{0}(G\times_{B}V_{-w_{0}(\lambda)})^{\ast}$ as $W_{\sqrt{p}Q}$-modules.
Proof.
It suffices to show the last assertion. By Lemma 2.9, Lemma 2.14 and Serre duality, we have
$H^{0}(G\times_{B}V_{\lambda})\simeq H^{l(w_{0})}(G\times_{B}V_{w_{0}\ast\lambda}(-\rho^{\prime}))\simeq H^{l(w_{0})}(G\times_{B}V_{-w_{0}(\lambda)}^{\ast}(-2\rho))\simeq H^{0}(G\times_{B}V_{-w_{0}(\lambda)})^{\ast}$
and thus the claim is proved, where $\rho^{\prime}$ in the second term is $\rho^{\vee}$ (case (1)) or $\rho$ (case (2)), respectively. ∎
3.1. The non-super case (1)
By applying Corollary 1.15 and Lemma 2.7, we obtain the following:
Corollary 3.3.
[45, Theorem 1.2, Lemma 4.21] For $\alpha\in P_{+}\cap Q$ and $\lambda\in\Lambda$ such that $(p\lambda_{\bullet}+\rho^{\vee},{}^{L}\theta)\leq p$,
(3.2) $\operatorname{ch}\mathcal{W}_{-\alpha+\lambda}=\sum_{\sigma\in W}(-1)^{l(\sigma)}\operatorname{ch}V_{\sigma\ast\lambda}^{h=\alpha+\lambda^{\bullet}-\sigma\uparrow\lambda}=\sum_{\sigma\in W}(-1)^{l(\sigma)}\tfrac{q^{\frac{1}{2p}|-p\sigma(\alpha+\lambda^{\bullet}+\rho)+p\lambda_{\bullet}+\rho^{\vee}|^{2}}}{\eta(q)^{r}}.$
Corollary 3.4.
[46, Section 3.2] For $m\geq{}^{L}h^{\vee}-1$ and $\lambda=\lambda^{\bullet}$, $W_{\sqrt{p}Q}$ and $W_{\sqrt{p}P}$ are simple as VOA and generalized VOA, respectively, and $W_{\sqrt{p}(Q-\lambda^{\bullet})}$ is simple as $W_{\sqrt{p}Q}$-module. Furthermore, for $m\geq{}^{L}h^{\vee}-1$ and $\beta\in P_{+}$, $\mathcal{W}_{-\beta}$ is simple as $\mathcal{W}_{0}$-module.
3.2. The super case (2)
In the same manner as above, we obtain the following:
Corollary 3.5.
For $\alpha\in P\cap Q$ and $\lambda\in\Lambda$ such that $(p\lambda_{\bullet}+\rho,{}^{L}\theta)\leq p$, we have
(3.3) $\operatorname{ch}\mathcal{W_{-\alpha+\lambda}}=\operatorname{ch}F\sum_{\sigma\in W}(-1)^{l(\sigma)}\operatorname{ch}V_{\sigma\ast\lambda}^{h=\alpha+\lambda^{\bullet}-\sigma\uparrow\lambda}=\operatorname{ch}F\sum_{\sigma\in W}(-1)^{l(\sigma)}\tfrac{q^{\Delta_{\sqrt{p}(-(\alpha+\lambda^{\bullet})+\sigma\ast\lambda_{\bullet})}}}{\eta(q)^{r}},$
(3.4) $\mathrm{sch}\mathcal{W}_{-\alpha+\lambda}=\mathrm{sch}F\sum_{\sigma\in W}(-1)^{l(\sigma)}\mathrm{sch}V_{\sigma\ast\lambda}^{h=\alpha+\lambda^{\bullet}-\sigma\uparrow\lambda}=\mathrm{sch}F\sum_{\sigma\in W}(-1)^{l(\sigma)+f(\sigma\circ(\alpha+\lambda^{\bullet}))}\tfrac{q^{\Delta_{\sqrt{p}(-(\alpha+\lambda^{\bullet})+\sigma\ast\lambda_{\bullet})}}}{\eta(q)^{r}},$
(3.5) $\operatorname{ch}\mathcal{S}_{R}\mathcal{W}_{-\alpha+\lambda}=q^{C+\tfrac{1}{16}}\operatorname{ch}\iota^{*}F\sum_{\sigma\in W}(-1)^{l(\sigma)}\tfrac{q^{\Delta_{\sqrt{p}(-(\alpha+\lambda^{\bullet})+\sigma\ast\lambda_{\bullet})}+(A\alpha_{r}+B\alpha_{r-1},-\sqrt{p}(\sigma\circ(\alpha+\lambda^{\bullet})-\lambda_{\bullet})}}{\eta(q)^{r}},$
where for any $\beta\in P$, define a function $f:P\mapsto\mathbb{Z}$ by $f(\beta)=\lfloor(\beta,\alpha_{r})\rfloor$, $A,B,C$ are defined in (2.23), and $\operatorname{sch}F$ and $\operatorname{ch}F$ are in Section 2.3. We denote $H^{0}(G\times_{B}\mathcal{S}_{R}V_{\lambda})=:\bigoplus_{\alpha\in P_{+}\cap Q}L_{\alpha+\lambda^{\bullet}}\otimes\mathcal{S}_{R}\mathcal{W}_{-\alpha+\lambda}$ by abuse of notation.
Corollary 3.6.
When $p\geq h^{\vee}-1$ and $\lambda=0$, $W_{\sqrt{p}Q}$ is simple as VOA and generalized VOA, respectively. Furthermore, for $\alpha\in P_{+}\cap Q$, $\mathcal{W}_{-\alpha}$ is simple as $\mathcal{W}_{0}$-module.
Remark 3.7.
For simply-laced $\mathfrak{g}$ (resp. rank 11 case), modularity of the characters were already studied in [11] (resp. [9, 8]). We expect similar results hold in our cases too.
4. Main results II
In Section 4, first we show that $\mathcal{W}_{0}$ is isomorphic to the corresponding principal W-algebra $\mathbf{W}^{k}(\mathfrak{g})$. It enables us to use another type character formula, i.e. Kazhdan–Lusztig type character formula (see [37, Theorem 1.1], [10, Theorem 7.7.1]) and the discussion in Section 2.2 (see the last two paragraphs). Using these results, we will attempt to extend the simplicity theorems (Corollary 3.4 and Corollary 3.6) to $\lambda\in\Lambda$ satisfying condition (strong) (or, equivalently, the conditions in Lemma 2.9). In the non-super case (1), the simplicity theorem will be proved in the same manner as [46]. On the other hand, the super case (2) has not been studied as much as the case (1), and we will prove the simplicity theorem under a particular assumption (exactness of ++-reduction). Since the flow of the discussion is the same as that of [46, Section 3.3], details are sometimes omitted.
4.1. Preliminary from W-algebra
Let us recall the notation in [34, Chapter 6] (see also [50, Chapter 1]) and [10, 5, 6]. For an affine Lie algebra $\hat{\mathfrak{g}}$, the Cartan subalgebra $\hat{\mathfrak{h}}$ and its dual $\hat{\mathfrak{h}}^{\ast}$ are decomposed as $\hat{\mathfrak{h}}=\mathfrak{h}\oplus(\mathbb{C}K+\mathbb{C}d)$ and $\hat{\mathfrak{h}}^{\ast}=\mathfrak{h}^{\ast}\oplus(\mathbb{C}\delta+\mathbb{C}\Lambda_{0})$, respectively. For $\hat{\mu}\in\hat{\mathfrak{h}}^{\ast}$, we have
$\hat{\mu}=\mu+\langle\hat{\mu},K\rangle\Lambda_{0}+(\mu,\Lambda_{0})\delta,\ \ (\mu\in\mathfrak{h}^{\ast}).$
[…]
$\sigma t_{\beta}(\hat{\mu})=\sigma(\mu+\langle\hat{\mu},K\rangle\beta)+\langle\hat{\mu},K\rangle\Lambda_{0}+((\hat{\mu},\Lambda_{0}-\beta)-\tfrac{1}{2}|\beta|^{2}\langle\hat{\mu},K\rangle)\delta.$
For $\hat{\rho}:=\rho+h^{\vee}\Lambda_{0}$, the circle $\hat{W}-$action on $\hat{\mathfrak{h}}^{\ast}$ is given by $w\circ\hat{\mu}:=w(\hat{\mu}+\hat{\rho})-\hat{\rho}$.
[…]
By applying the Drinfeld–Sokolov reduction $H^{0}_{\mathrm{DS}}(-)$ to $V^{k}(\mathfrak{g})$, the (universal) principal W-algebra $\mathbf{W}^{k}(\mathfrak{g})$ at level $k$ and its unique simple quotient $\mathbf{W}_{k}(\mathfrak{g})$ are obtained. Note that we have the Feigin–Frenkel duality
(4.1) $\mathbf{W}^{k}(\mathfrak{g})\simeq\mathbf{W}^{{}^{L}k}({}^{L}\mathfrak{g}),\ \ r^{\vee}(k+h^{\vee})({}^{L}k+{}^{L}h^{\vee})=1,\ \ \mathfrak{g}=X_{r}^{(1)},$
(4.2) $\mathbf{W}^{k}(\mathfrak{osp}(1|2r))\simeq\mathbf{W}^{{}^{L}k}(\mathfrak{osp}(1|2r)),\ \ 4(k+h^{\vee})({}^{L}k+h^{\vee})=1$
in our cases (see [22, 23], [28, Remark 6.5]). More generally, for $\check{\mu}\in\check{P}$ in the coweight lattice $\check{P}=Q^{\ast}$, we have the twisted Drinfeld–Sokolov reduction $H^{0}_{\mathrm{DS},\check{\mu}}(-)$ and the Arakawa–Frenkel module
$\mathbf{T}^{k+h^{\vee}}_{\beta,\check{\mu}}:=H^{0}_{\mathrm{DS},\check{\mu}}(\hat{V}_{\beta,k})$
over $\mathbf{W}^{k}(\mathfrak{g})$. For the Langlands dual $^{L}\mathfrak{g}$, we can define $\check{\mathbf{T}}^{{}^{L}k+{}^{L}h^{\vee}}_{\check{\mu},\beta}=H^{0}_{\mathrm{DS},\beta}(\hat{V}_{\beta,{}^{L}k})$ in the same manner. Then for $k\not\in Q$ or the case in [46], the Feigin–Frenkel duality $\mathbf{T}^{k+h^{\vee}}_{\beta,\check{\mu}}\simeq\check{\mathbf{T}}^{{}^{L}k+{}^{L}h^{\vee}}_{\check{\mu},\beta}$ under the identification $\mathbf{W}^{k}(\mathfrak{g})\simeq\mathbf{W}^{{}^{L}k}({}^{L}\mathfrak{g})$ is known to hold (see [5, 46]). Let $\chi_{\mu}:Z(\mathfrak{g})\rightarrow\mathbb{C}$ with $\mu\in\mathfrak{h}^{\ast}$ be the map in [6, (27)], where $Z(\mathfrak{g})$ is the center of the universal enveloping algebra of $U(\mathfrak{g})$. Then one can define the Verma module $\mathbf{M}_{k}(\chi_{\mu}):=U(\mathbf{W}^{k}(\mathfrak{g}))\otimes_{U(\mathbf{W}^{k}(\mathfrak{g})_{\geq 0})}\mathbb{C}_{\chi_{\mu}}$ over $\mathbf{W}^{k}(\mathfrak{g})$ with highest weight $\chi_{\mu}$ and its simple quotient $\mathbf{L}_{k}(\chi_{\mu})$ in the same manner above (in the Langlands dual case, denote $\check{\mathbf{M}}_{k}(\chi_{\mu})$ and $\check{\mathbf{L}}_{k}(\chi_{\mu})$, respectively).
4.2. The non-super case (1)
Let us consider our case (1). By the injectivity of the Miura map [28, Lemma 5.4] and the consideration in [26, Remark 4.1], we have
$\mathbf{W}^{k}(\mathfrak{g})\hookrightarrow\bigcap_{i\in I}\ker e^{-\frac{1}{{k+h^{\vee}}}b_{i}}_{(0)}|_{M(0)},$
where $b_{i}$ satisfies $(b_{i},b_{j})=(k+h^{\vee})(\alpha_{i},\alpha_{j})$. In particular, Fock module $M(\mu)$ is also a $\mathbf{W}^{k}(\mathfrak{g})$-module. By comparing the conformal vector (2.10) and [5, (3.22)], for $k=\frac{1}{p}-h^{\vee}$ and $b_{i}=-\frac{1}{\sqrt{p}}\alpha_{i}$, we have
[…]
On the other hands, by [5, Section 4.3], we have
(4.4) $\operatorname{ch}_{q}T^{m}_{p\lambda_{\bullet},\alpha+\lambda^{\bullet}}=\operatorname{ch}_{q}T^{\frac{1}{p}}_{\alpha+\lambda^{\bullet},p\lambda_{\bullet}}=\operatorname{ch}_{q}\mathcal{W}_{-\alpha+\lambda}$
for $\alpha\in P_{+}\cap Q$ and $\lambda\in\Lambda$ satisfying the strong condition Lemma 2.9(1). In particular, by applying $\alpha=\lambda=0$, we have $\operatorname{ch}_{q}\mathbf{W}^{m-{}^{L}h^{\vee}}({}^{L}\mathfrak{g})=\operatorname{ch}_{q}\mathbf{W}^{\frac{1}{p}-h^{\vee}}(\mathfrak{g})=\operatorname{ch}_{q}\mathcal{W}_{0}$, and thus the injection in (4.3) is isomorphic. By Corollary 3.4, $\mathcal{W}_{-\beta}$ is simple as $\mathbf{W}^{m-{}^{L}h^{\vee}}({}^{L}\mathfrak{g})$-module for any $\beta\in P_{+}$. In particular, for $\beta=0$, we have
(4.5) $\mathcal{W}_{0}\simeq\mathbf{W}^{m-{}^{L}h^{\vee}}({}^{L}\mathfrak{g})\simeq\mathbf{W}_{m-{}^{L}h^{\vee}}({}^{L}\mathfrak{g})\simeq\mathbf{W}^{\frac{1}{p}-h^{\vee}}(\mathfrak{g})\simeq\mathbf{W}_{\frac{1}{p}-h^{\vee}}(\mathfrak{g}).$
In the remaining of this subsection, we use
(4.6) $k=\tfrac{1}{p}-h^{\vee},\ \ \check{k}=m-{}^{L}h^{\vee}.$
On the other hand, in the same manner as [46, Lemma 3.14, 3.19], we have the following.
Lemma 4.1.
The top components of $M(\sqrt{p}(-\alpha+\lambda))$, $\mathcal{W}_{-\alpha+\lambda}$, $\mathbf{T}^{\frac{1}{p}}_{\alpha+\lambda^{\bullet},p\lambda_{\bullet}}$ and $\check{\mathbf{T}}^{m}_{p\lambda_{\bullet},\alpha+\lambda^{\bullet}}$ are $\mathbb{C}_{\chi_{\alpha-\lambda}}$ as $\operatorname{Zhu}(\mathbf{W}^{\frac{1}{p}-h^{\vee}}(\mathfrak{g}))$-module and $\mathbb{C}_{\chi_{p(-\alpha+\lambda)}}$ as $\operatorname{Zhu}(\mathbf{W}^{m-{}^{L}h^{\vee}}({}^{L}\mathfrak{g}))$-module, respectively. In particular, if the character of these modules coincides with that of the corresponding irreducible modules $\mathbf{L}_{k}({\chi_{\alpha-\lambda}})$ or $\check{\mathbf{L}}_{\check{k}}(\chi_{p(-\alpha+\lambda)})$, then they are isomorphic to $\mathbf{L}_{k}({\chi_{\alpha-\lambda}})$ or $\check{\mathbf{L}}_{\check{k}}(\chi_{p(-\alpha+\lambda)})$.
Our aim is to apply the discussion in Section 2.2 (see the last two paragraphs). First let us give a KL type decomposition in Lemma 2.6(3). Unless otherwise noted, all symbols (e.g., $Q$, $\alpha_{i}$,…) are the ones of $\mathfrak{g}$. Denote $h_{\mu}$ and $\check{h}_{\mu}$ the top conformal weight of $\mathbf{M}_{k}(\chi_{\mu})$ and $\check{\mathbf{M}}_{\check{k}}(\chi_{\mu})$, respectively (see [6, (30)]). By (2.6), for any $\beta\in(\sqrt{p}Q)^{\ast}$ (in other words, $\sqrt{p}\beta\in P_{{}^{L}\mathfrak{g}}$), we have
(4.7) $\Delta_{\beta}=h_{-\frac{1}{\sqrt{p}}\beta}=\check{h}_{\sqrt{p}\beta}.$
Let us take $\beta=\sqrt{p}(-\alpha+\lambda)$ ($\alpha\in P_{+}\cap Q$, $\lambda\in\Lambda$). Then the corresponding highest weight of $V^{\check{k}}({}^{L}\mathfrak{g})$ is
$-p(\alpha+\lambda^{\bullet}+\rho)+p\lambda_{\bullet}+\check{k}\Lambda_{0,{}^{L}\mathfrak{g}}.$
Let us recall the notation in [37]. In our case (i.e., $\widehat{{}^{L}\mathfrak{g}}$), we have
$\hat{W}_{{}^{L}\mathfrak{g}}=W_{{}^{L}\mathfrak{g}}\ltimes(Q_{{}^{L}\mathfrak{g}})^{\vee}=W\ltimes r^{\vee}Q,$
${\Delta}^{\mathrm{re}}_{+,\widehat{{}^{L}\mathfrak{g}}}=\{\alpha_{\pm}+n\delta\ |\ \alpha_{\pm}\in\Delta_{+,{}^{L}\mathfrak{g}},\ n\in\mathbb{Z}_{>0}\}\sqcup\Delta_{+,{}^{L}\mathfrak{g}}=\{\alpha_{\pm}^{\vee}+n\delta\ |\ \alpha_{\pm}\in\Delta_{+},\ n\in\mathbb{Z}_{>0}\}\sqcup\Delta_{+}^{\vee}$
$({\Delta}^{\mathrm{re}}_{+,\widehat{{}^{L}\mathfrak{g}}})^{\vee}=\{r^{\vee}\alpha_{\pm}^{l}+r^{\vee}n\delta\ |\ \alpha_{\pm}^{l}\in\Delta_{\pm}^{l},\ n\in\mathbb{Z}_{>0}\}\sqcup\{r^{\vee}\alpha_{\pm}^{s}+n\delta\ |\ \alpha_{\pm}^{s}\in\Delta_{\pm}^{s},\ n\in\mathbb{Z}_{>0}\}\sqcup r^{\vee}\Delta_{+},$
$\begin{aligned} {\mathcal{C}}^{+}_{\widehat{{}^{L}\mathfrak{g}}}&=\{\nu\ |\ (\gamma^{\vee},\nu+\hat{\rho}_{{}^{L}\mathfrak{g}})_{{}^{L}Q}\geq 0\ \text{for any $\gamma\in\Delta^{\mathrm{re}}_{+,\widehat{{}^{L}\mathfrak{g}}}$}\}\\ &=\{\nu\ |\ (\gamma^{\vee},\nu+\rho^{\vee}+{}^{L}h^{\vee}\Lambda_{0,{}^{L}\mathfrak{g}})_{{}^{L}Q}\geq 0\ \text{for any $\gamma^{\vee}\in(\Delta^{\mathrm{re}}_{+,\widehat{{}^{L}\mathfrak{g}}})^{\vee}$}\}.\end{aligned}$
By [37, Lemma 2.10], there exists unique $\sigma t_{r^{\vee}\beta}\in\hat{W}_{{}^{L}\mathfrak{g}}=W\ltimes r^{\vee}Q$ such that
$\sigma t_{r^{\vee}\beta}\circ_{\widehat{{}^{L}\mathfrak{g}}}(-p(\alpha+\lambda^{\bullet}+\rho)+p\lambda_{\bullet}+\check{k}\Lambda_{0,{}^{L}\mathfrak{g}})\in{\mathcal{C}}^{+}_{\widehat{{}^{L}\mathfrak{g}}}$.
In other words, we have
$0\leq(\sigma^{-1}(\Delta_{+}^{l}),p(\beta-(\alpha+\lambda^{\bullet}+\rho))+p\lambda_{\bullet}+\rho^{\vee})_{Q}\leq r^{\vee}m,$
$\land\$$0\leq(\sigma^{-1}(\Delta_{+}^{s}),p(\beta-(\alpha+\lambda^{\bullet}+\rho))+p\lambda_{\bullet}+\rho^{\vee})_{Q}\leq m.$
We show that if $\lambda$ satisfies the condition Lemma 2.9(1), i.e. $(p\lambda_{\bullet}+\rho^{\vee},\theta_{s})\leq m$, such $\sigma t_{r^{\vee}\beta}$ is independent of the choice of $\lambda_{\bullet}$. Let us consider the case $\mathfrak{g}=B_{r}$ (other cases are similar). Set $\beta-(\alpha+\lambda^{\bullet}+\rho)=\sum_{i\in I}n_{i}\varpi_{i}$ ($n_{i}\in\mathbb{Z}$). For $\alpha_{+}\in\Delta_{+}$, if there exists, denote $a_{\pm\alpha_{+}}\in\Delta_{+}$ such that $\sigma^{-1}(a_{\pm\alpha_{+}})=\pm\alpha_{+}$. For any $\alpha_{+}\in\Delta_{+}$, one of $a_{\alpha_{+}}$ or $a_{-\alpha_{+}}$ always exists. Set $I_{\pm}=\{i\ |\ \exists a_{\pm\alpha_{i}}\}$. Clearly, we have $I=I_{+}\sqcup I_{-}$. For $i\in I_{+}$, we have
$0\leq(\alpha_{i},p\sum_{i\in I}n_{i}\varpi_{i}+p\lambda_{\bullet}+\rho^{\vee})=pn_{i}(\alpha_{i},\varpi_{i})+\lambda_{i}\leq\begin{cases}r^{\vee}m&(\alpha_{i}\in\Delta_{+}^{l}),\\ m&(\alpha_{i}\in\Delta_{+}^{s}),\end{cases}$
and thus $n_{i}=0$ for $i\in I_{+}$. Similarly, we have $n_{i}=-1$ for $i\in I_{-}$, and thus $\sum_{i\in I}n_{i}\varpi_{i}=-\sum_{i\in I_{-}}\varpi_{i}$. Next, let us consider $a_{\pm\theta_{s}}$. If $a_{\theta_{s}}$ exists, then we have $I_{-}=\phi$. On the other hand, if $a_{-\theta_{s}}$ exists, then we have $I_{-}=\{r\}$. Each of these cases corresponds to the case $\alpha+\lambda^{\bullet}+\rho\in Q$ or not, which is independent of the choice of $\lambda_{\bullet}$. Therefore, we can use the notation $y_{\alpha,\lambda^{\bullet}}\in\hat{W}$ for the inverse of the unique element $\sigma t_{\beta}$ in above discussion. By [37, Theorem 1.1], we obtain the following:
Lemma 4.2.
For $\alpha\in P_{+}\cap Q$ and $\lambda\in\Lambda$ in (2.12) s.t. $(p\lambda_{\bullet}+\rho^{\vee},{}^{L}\theta)\leq p$, we have
$\operatorname{ch}\hat{L}_{\check{k}}(-p(\alpha+\lambda^{\bullet}+\rho)+p\lambda_{\bullet}+\check{k}\Lambda_{0,{}^{L}\mathfrak{g}})_{{}^{L}\mathfrak{g}}=\sum_{y\geq y_{\alpha,\lambda^{\bullet}}}a_{y,y_{\alpha,\lambda^{\bullet}}}\operatorname{ch}\hat{M}_{\check{k}}(y\circ_{\widehat{{}^{L}\mathfrak{g}}}\mu_{\lambda})_{{}^{L}\mathfrak{g}},$
where $a_{y,w}=Q_{y,w}(1)$ for the inverse Kazhdan-Lusztig polynomial $Q_{y,w}(q)$ and
(4.8) $\mu_{\lambda}=y_{\alpha,\lambda^{\bullet}}^{-1}\circ_{\widehat{{}^{L}\mathfrak{g}}}(-p(\alpha+\lambda^{\bullet}+\rho)+p\lambda_{\bullet}+\check{k}\Lambda_{0,{}^{L}\mathfrak{g}}).$
Furthermore, by applying the (exact) ++-reduction functor (see [10, Theorem 7.7.1]), we have
$\operatorname{ch}\check{\mathbf{L}}_{\check{k}}(\chi_{p(-\alpha+\lambda)})=\sum_{y\geq y_{\alpha,\lambda^{\bullet}}}a_{y,y_{\alpha,\lambda^{\bullet}}}\operatorname{ch}\check{\mathbf{M}}_{\check{k}}(\chi_{y\circ_{\widehat{{}^{L}\mathfrak{g}}}\mu_{\lambda}+p\rho-\check{k}\Lambda_{0,{}^{L}\mathfrak{g}}}).$
In particular, the condition Lemma 2.6(3) holds in this case.
Let us apply the discussion in Section 2.2 to our case. In Lemma 2.6, set
$\beta=\alpha+\lambda^{\bullet},\ \ (\lambda_{0},\lambda_{1})=(0,\lambda_{\bullet}),\ \ a_{y,\beta}=Q_{y,y_{\alpha,\lambda^{\bullet}}}(1),$
[…]
${\mathbb{M}}(y,\mu_{\lambda_{i}})=\check{\mathbf{M}}_{\check{k}}(\chi_{y\circ_{\widehat{{}^{L}\mathfrak{g}}}\mu_{\lambda}+p\rho-\check{k}\Lambda_{0,{}^{L}\mathfrak{g}}}),\ \ {\mathbb{L}}(-\beta+\lambda_{i})=\check{\mathbf{L}}_{\check{k}}(\chi_{p(-\alpha+\lambda)}),$
respectively. Then the conditions Lemma 2.6(3),(4) are already checked. On the other hand, the condition Lemma 2.6(1),(2) follows from (4.7) and [6, (29)]. Then by Lemma 2.6 and Lemma 4.1, $\mathcal{W}_{-\alpha+\lambda}$ is simple as $\mathcal{W}_{0}$-module. In particular, each $\mathcal{W}_{-\alpha+\lambda}$ is a CFT-type $\mathcal{W}_{0}$-module. Finally, by applying Lemma 2.8, the simplicity of $H^{0}(G\times_{B}V_{\lambda})$ also holds. In conclusion, we have the following:
Theorem 4.3.
Let us consider the non-super case (1) and $H^{0}(G\times_{B}V_{\lambda})\simeq\bigoplus_{\alpha\in P_{+}\cap Q}L_{\alpha+\lambda^{\bullet}}\otimes\mathcal{W}_{-\alpha+\lambda}$. Then, for (4.6) and $m\geq{}^{L}h^{\vee}-1$, we have the VOA-isomorphism
$\mathcal{W}_{0}\simeq\mathbf{W}^{\check{k}}({}^{L}\mathfrak{g})\simeq\mathbf{W}_{\check{k}}({}^{L}\mathfrak{g})\simeq\mathbf{W}^{k}(\mathfrak{g})\simeq\mathbf{W}_{k}(\mathfrak{g})$
and each $\mathcal{W}_{-\alpha+\lambda^{\bullet}}$ is simple as $\mathcal{W}_{0}$-module. More generally, for $\lambda\in\Lambda$ such that $(p\lambda_{\bullet}+\rho^{\vee},{}^{L}\theta)\leq p$,
$\mathcal{W}_{-\alpha+\lambda}$ $\simeq\check{\mathbf{L}}_{\check{k}}(\chi_{p(-\alpha+\lambda)})\simeq\check{\mathbf{T}}^{m}_{p\lambda_{\bullet},\alpha+\lambda^{\bullet}}\ \ \text{(as $\mathbf{W}^{\check{k}}({}^{L}\mathfrak{g})$-modules)}$
$\simeq\mathbf{L}_{k}(\chi_{\alpha-\lambda})\simeq\mathbf{T}^{\frac{1}{p}}_{\alpha+\lambda^{\bullet},p\lambda_{\bullet}}\ \ \text{(as $\mathbf{W}^{k}(\mathfrak{g})$-modules)},$
and $H^{0}(G\times_{B}V_{\lambda})$ is simple as $H^{0}(G\times_{B}V_{0})$-module, respectively.
Remark 4.4.
It is expected that the simple modules of $H^{0}(G\times_{B}V_{0})$ are classified by $\Lambda$. Constructing simple modules for general $\lambda\in\Lambda$ seems to require the development of not only our geometric approach, but also the W-algebraic analysis of the Verma module $)\mathbf{M}_{k}(\chi_{\mu})=H^{0}_{\mathrm{DS},\check{\mu}}(\hat{M}(\hat{\lambda}$)). The latter is currently being pursued by T. Arakawa and the first author to investigate the tensor category structure of W-algebras at generic levels.
4.3. The super case (2)
In this subsection, set
(4.9) ${k}=\tfrac{1}{2(2m-1)}-(r+\tfrac{1}{2}),\ \ \check{k}=m-r-1,\ \ h^{\vee}=h^{\vee}_{\mathfrak{osp}(1|2r)}=r+\tfrac{1}{2}.$
Then $4(k+h^{\vee})(\check{k}+h^{\vee})=1$ and by [28, Theorem 6.3] and Corollary 3.6, we have
(4.10) $\mathcal{W}_{0}\simeq\mathbf{W}^{k}(\mathfrak{osp}(1|2r))\simeq\mathbf{W}_{k}(\mathfrak{osp}(1|2r))\simeq\mathbf{W}^{\check{k}}(\mathfrak{osp}(1|2r))\simeq\mathbf{W}_{\check{k}}(\mathfrak{osp}(1|2r))$
and each $\mathcal{W}_{-\alpha}$ ($\alpha\in P_{+}\cap Q$) is simple as $\mathcal{W}_{0}$-module. In the same manner as above, we have the following:
Lemma 4.5.
The top component of $M(-\sqrt{p}(\alpha+\lambda))$ and $\mathcal{W}_{-\alpha+\lambda}$ are $\mathbb{C}_{\chi_{\alpha-\lambda}}$ as $\operatorname{Zhu}(\mathbf{W}_{{k}}(\mathfrak{osp}(1|2r))$-modules and $\mathbb{C}_{\chi_{p(-\alpha+\lambda)}}$ as $\operatorname{Zhu}(\mathbf{W}_{\check{k}}(\mathfrak{osp}(1|2r))$-modules, respectively. In particular, if the character of these modules coincides with that of $\mathbf{L}_{{k}}(\chi_{-\alpha+\lambda})$ or $\mathbf{L}_{\check{k}}(\chi_{p(-\alpha+\lambda)})$, then they are isomorphic to $\mathbf{L}_{{k}}(\chi_{-\alpha+\lambda})$ or $\mathbf{L}_{\check{k}}(\chi_{p(-\alpha+\lambda)})$.
We aim to apply the discussion in Section 2.2 (see the last two paragraphs). First let us consider Lemma 2.6 in our case ($\lambda_{0}=0$, $\lambda_{1}=\lambda$ such that $\lambda^{\bullet}=0$ and $(p\lambda_{\bullet}+\rho,{}^{L}\theta)\leq p$). The condition Lemma 2.6(4) is already checked in the above discussion. By [36] and the results in Section 2.3, we have
$\operatorname{ch}_{q}\mathbf{M}_{\check{k}}(\chi_{\mu})=\tfrac{\operatorname{ch}_{q}F}{\eta(q)^{r}}{q^{\frac{1}{2p}|\mu-p\rho+\frac{1}{p}\rho|^{2}}},\ \ \text{in particular, $\operatorname{ch}_{q}\mathbf{M}_{\check{k}}(\chi_{p(-\alpha+\lambda)})=\operatorname{ch}_{q}V_{\lambda}^{h=\sigma\circ(\alpha+\lambda^{\bullet})}$ }$
and the condition Lemma 2.6(1),(2) are satisfied.
Lastly, we want to check the Lemma 2.6(3). However, to the authors’ knowledge, the KL type decomposition as [37, 10] has not been proven in our case $\mathfrak{osp}(1|2r)$. In this paper, we will prove the KL type decomposition (and Lemma 2.6(3)) in the case of twisted affine Lie algebra of type $A^{(2)}_{2r}$, and prove the simplicity theorem for super-case (2) under the assumption that the characters in these two types are the same under the coordinate change. Let us recall the data in $\hat{\mathfrak{g}}^{\prime}=A_{2r}^{(2)}$: [^11] Note that here we use the second description in [12]. On the other hand, [34, 50] uses the first description of [12]. The classical part is different for these two descriptions: For the first one (resp. second one), it is type $C_{n}$ (resp. type $B_{n}$).
$\hat{W}_{\hat{\mathfrak{g}}^{\prime}}=W\ltimes Q,\ \ h^{\vee}_{A_{2r}^{(2)}}=2r+1=2h^{\vee}_{\mathfrak{osp}(1|2r)},$
${\Delta}^{\mathrm{re}}_{+,\hat{\mathfrak{g}}^{\prime}}=\{\alpha_{\pm}+n\delta\ |\ \alpha_{\pm}\in\Delta_{\pm},\ n\in\mathbb{Z}_{>0}\}\sqcup\{2\alpha_{\pm}^{s}+(2n+1)\delta\ |\ \alpha_{\pm}^{s}\in\Delta_{\pm}^{s},\ n\in\mathbb{Z}_{\geq 0}\}\sqcup\Delta_{+},$
${\mathcal{C}}^{+}_{\hat{\mathfrak{g}}^{\prime}}=\{\mu\in\hat{\mathfrak{h}}^{\ast}\ |\ (\gamma^{\vee},\mu+\hat{\rho}_{A_{2r}^{(2)}})\in\mathbb{Z}_{\geq 0}\ \text{for any $\gamma\in{\Delta}^{\mathrm{re}}_{+,\hat{\mathfrak{g}}^{\prime}}$}\},\ \ \hat{\rho}_{A_{2r}^{(2)}}=\rho+h^{\vee}_{A_{2r}^{(2)}}\Lambda_{0}^{c},$
where all classical terms are of type $B_{n}$ (see [34, Section 6]), and $\Lambda_{0}^{c}=\frac{1}{2}\Lambda_{0}$. By [37, Lemma 2.10], there exists unique $\sigma t_{\beta}\in\hat{W}_{\hat{\mathfrak{g}}^{\prime}}=W\ltimes Q$ such that
$\sigma t_{\beta}\circ_{\hat{\mathfrak{g}}^{\prime}}(-p(\alpha+\lambda^{\bullet}+\rho^{\vee})+p\lambda_{\bullet}+\check{k}\Lambda_{0})=\sigma(p(\beta-\alpha-\lambda^{\bullet}-\rho^{\vee})+p\lambda_{\bullet}+\rho)+p\Lambda_{0}^{c}-\hat{\rho}_{A_{2r}^{(2)}}\in{\mathcal{C}}^{+}_{\hat{\mathfrak{g}}^{\prime}}$.
In other words, we have
$0\leq(p(\beta-\alpha-\lambda^{\bullet}-\rho^{\vee})+p\lambda_{\bullet}+\rho,\sigma^{-1}(\alpha_{+}^{\vee}))\leq p\ \ \text{for any $\alpha_{+}\in\Delta_{+}$}.$
In the same manner as Section 4.2, we can determine such a ($\lambda_{\bullet}$-independent) $\sigma t_{\beta}$ explicitly as follows:
$\begin{cases}(\sigma,\beta)=(\operatorname{id},\alpha+\rho^{\vee}),\ \ \mu_{\lambda}=p\lambda+\check{k}\Lambda_{0}&(\lambda^{\bullet}=0)\\ (\sigma,\beta)=(\sigma_{r},\alpha+\rho^{\vee}),\ \ \mu_{\lambda}=\sigma_{r}\circ_{\hat{\mathfrak{g}}^{\prime}}(p\lambda+\check{k}\Lambda_{0})&(\lambda^{\bullet}=\varpi_{r})\end{cases}$
By combining it with [37, Theorem 1.1], we obtain the KL type decomposition in $A_{2r}^{(2)}$ case.
Lemma 4.6.
For $\alpha\in P_{+}\cap Q$ and $\lambda\in\Lambda$ in (2.16) s.t. $(p\lambda_{\bullet}+\rho,{}^{L}\theta)\leq p$, we have
$\operatorname{ch}\hat{L}(-p(\alpha+\lambda^{\bullet}+\rho^{\vee})+p\lambda_{\bullet}+\check{k}\Lambda_{0})_{\hat{\mathfrak{g}}^{\prime}}=\sum_{y\geq y_{\alpha,\lambda^{\bullet}}}a_{y,y_{\alpha,\lambda^{\bullet}}}\operatorname{ch}\hat{M}(y\circ_{\hat{\mathfrak{g}}^{\prime}}\mu_{\lambda})_{\hat{\mathfrak{g}}^{\prime}},$
where $a_{y,w}=Q_{y,w}(1)$ for the inverse Kazhdan-Lusztig polynomial $Q_{y,w}(q)$ and
$y_{\alpha,\lambda^{\bullet}}=\begin{cases}t_{-\alpha-\rho^{\vee}}&(\lambda^{\bullet}=0),\\ t_{-\alpha-\rho^{\vee}}\sigma_{r}&(\lambda^{\bullet}=\varpi_{r}),\end{cases}\ \ \mu_{\lambda}=\begin{cases}p\lambda+\check{k}\Lambda_{0}&(\lambda^{\bullet}=0),\\ \sigma_{r}\circ_{\hat{\mathfrak{g}}^{\prime}}(p\lambda+\check{k}\Lambda_{0})&(\lambda^{\bullet}=\varpi_{r}).\end{cases}$
Let us relate type $A_{2r}^{(2)}$ and $\mathfrak{osp}(1|2r)$. Because the second description of $A_{2n}^{(2)}$ in [12] has the same Dynkin diagram as the one of type $B(0,r)=\mathfrak{osp}(1|2r)$, replacing a black node with a white node, these two types are closely related. In particular, according to [35, Section 9.5], the character formulas for the modules of affine Lie (super)algebras of type $A_{2r}^{(2)}$ and of type $B(0,r)$ are the same up to coordinate change. In fact, choosing the coordinate of $h\in\hat{\mathfrak{h}}$ of twisted affine Kac-Moody Lie algebra of type $A_{2r}^{(2)}$:
$h:=2\pi i(-\tau\Lambda_{0}^{c}+z+u\delta).$
Then by proper coordinate transformation $t$, one can obtain that
(4.11) $\operatorname{ch}[\hat{M}(A_{2r}^{(2)})](h)=\operatorname{ch}[\hat{M}(B(0,r)](t(h)).$
Meanwhile, for type $A_{2n}^{(2)}$, the KL type decomposition was already proved in Lemma 4.6 above. Therefore, we also have the same KL type decomposition in $\mathfrak{osp}(1|2r)$-side.
To apply Lemma 2.6(3) to our super case (2), one must move from affine to the principal $W$-algebra. We denote the category $\mathcal{O}$ for $L_{\kappa}(\mathfrak{osp}(1|2r))$ by $\mathcal{O}_{\kappa}$ ($\kappa\in\mathbb{C}$). Let $\mathcal{O}_{\kappa}^{[\lambda]}$ be the full subcategory of $\mathcal{O}_{\kappa}$ whose objects have their local composition factors isomorphic to $\hat{L}_{\kappa}(w\circ\lambda)$. The exactness of the ++-reduction $H^{0}_{\mathrm{DS},+}(\cdot)$ is already proved in [10] (and thus the KL type character formula of $\mathbf{W}^{\kappa}(\mathfrak{g})$-module is derived from [37, Theorem 1.1]), but in our case it has not known yet to the authors’ knowledge.
Conjecture 4.7.
For $\check{k}=m-r-1$ and $\lambda\in\Lambda$ $(p\lambda_{\bullet}+\rho,{}^{L}\theta)\leq p$, the ++-reduction functor $H^{0}_{\mathrm{DS},+}(\cdot)$ defines an exact functor from $\mathcal{O}_{\check{k}}^{[\lambda]}$ to $\mathbf{W}_{\check{k}}(\mathfrak{osp}(1|2r))\text{-$\mathrm{mod}$}$.
If we assume the above conjecture, then we have the KL type decomposition in our case $\mathbf{W}^{\check{k}}(\mathfrak{osp}(1|2r))$ and the last condition Lemma 2.6(3) is also satisfied. The remaining discussion is the same as Section 4.2, and the conclusion is as follows:
Theorem 4.8.
Let us consider the super case (2), and set $H^{0}(G\times_{B}V_{\lambda})\simeq\bigoplus_{\alpha\in P_{+}\cap Q}L_{\alpha+\lambda^{\bullet}}\otimes\mathcal{W}_{-\alpha+\lambda}$. Then, for (4.9) and $m\geq r$, we have the VOSA-isomorphism
$\mathcal{W}_{0}\simeq\mathbf{W}^{\check{k}}(\mathfrak{osp}(1|2r))\simeq\mathbf{W}_{\check{k}}(\mathfrak{osp}(1|2r))\simeq\mathbf{W}^{k}(\mathfrak{osp}(1|2r))\simeq\mathbf{W}_{k}(\mathfrak{osp}(1|2r))$
for $\check{k}=m-r-1$, $k=\tfrac{1}{2(2m-1)}-(r+\tfrac{1}{2})$, and each $\mathcal{W}_{-\alpha}$ is simple as $\mathcal{W}_{0}$-module ($\alpha\in P_{+}\cap Q$). Furthermore, under the assumption of Conjecture 4.7, for any $\alpha\in P_{+}\cap Q$ and $\lambda\in\Lambda$ such that $\lambda^{\bullet}=0$ and $(p\lambda_{\bullet}+\rho^{\vee},{}^{L}\theta)\leq p$,
$\mathcal{W}_{-\alpha+\lambda}$ $\simeq{\mathbf{L}}_{\check{k}}(\chi_{p(-\alpha+\lambda)})\ \ \text{(as $\mathbf{W}^{\check{k}}(\mathfrak{osp}(1|2r))$-modules)}$
$\simeq\mathbf{L}_{k}(\chi_{\alpha-\lambda})\ \ \text{(as $\mathbf{W}^{k}(\mathfrak{osp}(1|2r))$-modules)},$
and $H^{0}(G\times_{B}V_{\lambda})$ is simple as $H^{0}(G\times_{B}V_{0})$-module, respectively. Finally, if one can check the simplicity for one $\lambda\in\Lambda$ such that $\lambda^{\bullet}=\varpi_{r}$ and $(p\lambda_{\bullet}+\rho,{}^{L}\theta)\leq p$, then the simplicity theorem for $\mathcal{W}_{-\alpha+\lambda}$ and $H^{0}(G\times_{B}V_{\lambda})$ above holds for any $\lambda\in\Lambda$ satisfying the same condition.